### v2

# EvolveNav: Self-Improving Embodied Reasoning for LLM-Based Vision-Language Navigation

Abstract
Building Vision-Language Navigation (VLN) agents which can navigate following natural language instructions is a long-standing goal in human-robot interaction applications. Recent studies have revealed the potential of training open-source Large Language Models (LLMs) to unleash LLMs’ reasoning ability for improving navigation, and simultaneously mitigate the domain gap between LLMs’ training corpus and the VLN task. However, these approaches primarily adopt direct input-output mapping paradigms, causing the mapping learning difficult and the navigational decisions unexplainable. Chain-of-Thought (CoT) training is a promising way to improve both navigational decision accuracy and interpretability, while the complexity of the navigation task makes the perfect CoT labels unavailable and may lead to overfitting through pure CoT supervised fine-tuning. In this paper, we propose a novel sElf-improving embodied reasoning framework for boosting LLM-based vision-language Navigation, dubbed EvolveNav. Our EvolveNav consists of two stages: (1) Formalized CoT Supervised Fine-Tuning, where we train the model with formalized CoT labels to both activate the model’s navigational reasoning capabilities and increase the reasoning speed; (2) Self-Reflective Post-Training, where the model is iteratively trained with its own reasoning outputs as self-enriched CoT labels to enhance the supervision diversity. A self-reflective auxiliary task is also introduced to encourage learning correct reasoning patterns by contrasting with wrong ones. Experimental results on the popular VLN benchmarks demonstrate the superiority of EvolveNav over previous LLM-based VLN approaches. Code is available at: https://github.com/expectorlin/EvolveNav.
1 Introduction
Vision-Language Navigation (VLN) has received significant research interest within the Embodied AI community, due to its practicality and flexibility in enabling human-robot interaction in real-world robotic applications. In VLN tasks, an embodied agent needs to follow natural language instructions to navigate through complex visual environments to reach the target position. Early works improve VLN performance by designing dedicated model architectures wang2019reinforced ; ma2019self ; deng2020evolving ; qi2020Object , introducing powerful learning paradigms zhu2020vision ; li2019robust ; huang2019transferable , and developing useful data augmentation techniques tan2019learning ; fried2018speaker ; liu2021vision ; Fu2019CounterfactualVN . Subsequently, pretraining-based VLN approaches have been widely proposed to improve the cross-modal alignment ability and decision accuracy of navigation agents hong2021vln ; Chen2021HistoryAM ; Qi2021TheRT ; Chen2022ThinkGA ; Qiao2022HOPHA . Nevertheless, constrained by the limited scale of pretraining and VLN in-domain data, these approaches cannot learn navigational reasoning knowledge sufficiently, and therefore still struggle to handle various unseen navigation scenarios.
With the rapid progress of large language models (LLMs) brown2020language ; touvron2023llama ; touvron2023llama2 , emerging works have introduced LLMs to address embodied tasks considering LLMs’ rich real-world common sense and powerful reasoning ability Ahn2022DoAI ; Huang2022InnerME ; driess2023palme . Some recent works have attempted to build LLM-based VLN models in a zero-shot or trainable manner  chen2024mapgpt ; zhou2023navgpt ; long2023discuss ; hao2020towards ; zhang2024navid . The zero-shot approaches, such as NavGPT zhou2023navgpt and MapGPT chen2024mapgpt , resort to closed-source LLMs OpenAI_2023 to generate navigational reasoning and decision for different navigation timesteps. To alleviate the high cost of frequently querying closed-source LLMs for sequential decision making, the trainable methods collect in-domain data to train open-source LLMs vicuna to build the navigation agent. However, they typically map directly from navigational inputs to decisions without explicit intermediate reasoning steps, leading to decision uninterpretability and may also limit performance (see Figure 1(a)).
Figure 1: Comparison of different chain-of-thought (CoT) training paradigms. (a) Direct Mapping Learning maps the navigation inputs to actions straightforwardly. (b) Formalized CoT Learning and (c) Free-form CoT Learning generate formalized and free-form reasoning, respectively, under the training with fixed CoT labels. (d) Different from the above paradigms, our Self-Improving CoT Learning framework utilizes the model’s own reasoning outputs as self-enriched CoT labels and learn the reasoning in a self-reflective way during CoT training to fulfill generalizable and adaptable reasoning. Red and green fonts represent wrong and correct reasoning outputs, respectively. R+ and R- represent positive and negative reasoning samples, respectively.
Recent studies have revealed the effectiveness of embodied chain-of-thoughts (CoT) training in enhancing both the decision accuracy and interpretability for embodied tasks zawalski2024robotic ; zhao2025cot ; lin2024navcot . Most of these approaches employ the supervised fine-tuning paradigm for embodied CoT training. However, introducing CoT supervised fine-tuning for training VLN models is challenging. Firstly, due to the complexity and uncertainty of the navigation task, there can be multiple cues for deciding the correct navigation action, i.e., there can be multiple “correct” CoT supervisions to guide navigation for a specific timestep. This leads to a hard collection process of perfect navigation CoT labels. Secondly, pure CoT supervised fine-tuning using fixed CoT labels may cause overfitting to certain reasoning patterns and harm the generalization to diverse unseen scenarios.
In this paper, we propose a sElf-improving embodied reasoning paradigm for enhancing LLM-based vision-language Navigation, called EvolveNav, to fulfill generalizable and adaptable navigational reasoning under different scenes. Our EvolveNav comprises two stages: 1) Formalized CoT Supervised Fine-Tuning and 2) Self-Reflective Post-Training. The Stage 1 training of Formalized CoT Supervised Fine-Tuning aims to first activate the model’s potential reasoning capabilities, where we ask the model to produce explicit chain-of-thought navigational reasoning by predicting the landmarks needed to locate with the corresponding direction for deciding the navigation actions. To alleviate redundant reasoning and increase the inference speed, we conduct the CoT supervised fine-tuning using formalized CoT labels, which are collected by filling the landmark and direction information into concise label templates. To mitigate the overfitting to pre-constructed CoT labels and enable self-improving reasoning for enhancing generalization, we further conduct Self-Reflective Post-Training for Stage 2 training. Concretely, we design a self-enriched CoT label enhancement strategy, where we train the model with its iteratively produced correct reasoning outputs to enrich the CoT supervision. We also introduce a self-reflective auxiliary task, where the model needs to discriminate between positive and negative navigational reasoning to learn correct reasoning patterns better. As shown in Figure 1, in contrast to CoT supervised fine-tuning using fixed CoT labels (i.e., Figure 1(b) Formalized CoT Learning and (c) Free-form CoT Learning), our EvolveNav (Figure 1(d)) can generate embodied CoT in a self-refining manner during training to mitigate the overfitting. Additionally, through training with formalized CoT labels, our EvolveNav can significantly reduce uninformative navigational reasoning to promote the reasoning speed compared with using free-form CoT labels for training (as in Figure 1 (c)).
We conduct experiments on two public VLN benchmarks, R2R anderson2018vision and CVDN thomason2019vision . Experimental results show that EvolveNav substantially outperforms previous LLM-based VLN approaches on unseen scenarios, demonstrating the effectiveness of our self-improving embodied reasoning paradigm in promoting navigation decision accuracy and generalization. We carefully conduct ablation experiments to explore how to design streamlined CoTs that can provide interpretability while boosting navigation performance. Visualization also insightfully reveals the reasonability of our design for CoT labels in enhancing decision interpretability and improving navigational reasoning.
To summarize, the main contributions of this paper are:
  • We propose a novel self-improving embodied reasoning paradigm, EvolveNav, for enhancing the LLM-based vision-and-language navigation, which fulfills generalizable and adaptable navigational reasoning under unseen scenes.
  • We construct formalized CoT labels for conducting supervised fine-tuning, which effectively activates the agent’s navigational reasoning ability and promotes the reasoning speed. We introduce a self-enriched CoT label enhancement strategy and a self-reflective auxiliary task to enable learning correct reasoning patterns in a self-refining manner to mitigate overfitting.
  • Experimental results show the superiority of EvolveNav over other LLM-based approaches on two standard VLN benchmarks R2R and CVDN. Our EvolveNav can improve the generalization of both navigational reasoning and decision-making, providing meaningful insights for designing further embodied reasoning paradigms.
2 Related Work
Vision-Language Navigation (VLN). Vision-Language Navigation (VLN) has attracted intensive research interest in recent years. Various VLN benchmarks have been proposed to evaluate agents’ ability for navigational reasoning and instruction following anderson2018vision ; jain2019stay ; ku2020room ; thomason2019vision ; qi2020reverie ; zhu2021soon . Previous approaches employ non-pretraining-based wang2019reinforced ; ma2019self ; deng2020evolving ; zhu2020vision ; tan2019learning ; fried2018speaker or pretraining-based paradigms hong2021vln ; Chen2021HistoryAM ; Chen2022ThinkGA ; Qiao2022HOPHA ; Guhur2021AirbertIP ; an2022bevbert ; wang2023scaling for tackling the above VLN tasks. However, these approaches cannot generalize well to diverse unseen scenarios that require rich real-world commonsense, and the navigation decisions also lack explainability. Some recent works have introduced LLMs to assist the VLN task, by either eliciting the useful navigation knowledge stored in LLMs lin2024correctable ; zhou2024navgpt ; qiao2023march or employing the LLM as the navigation backbone for action decision zhou2023navgpt ; zheng2024towards ; zhang2024navid ; lin2024navcot ; chen2024mapgpt . Our work lies in the latter. Different from previous approaches, we propose a new VLN framework in this work, where the LLM-based navigation backbone iteratively generates intermediate reasoning steps in a self-improving manner during training to guide navigational decisions. As a result, both the reasoning ability and decision interpretability of the navigation model can be significantly enhanced.
LLMs as Embodied Agents. Recent research have revealed the giant potential of utilizing Large Language Models (LLMs) as embodied agents to complete the robotic navigation and manipulation tasks, benefiting from the outstanding ability of planning, reasoning, and reflection of LLMs song2023llmplanner ; Ahn2022DoAI ; Huang2022InnerME ; Yao2022ReActSR ; shahlm ; schumann-2023-velma ; wang2023voyager ; huang2023voxposer ; rajvanshi2024saynav . For example, LM-Nav shahlm introduces the LLM to parse the long navigation instruction into sequential landmarks for facilitating the navigational planning. Voxposer huang2023voxposer introduces LLMs for code writing and combines them with the Vision-Language models (VLMs) to compose 3D value maps for robotic manipulation.
There are typically two branches of works where the LLMs act as embodied agents for tackling the VLN task. In the first branch, closed-source LLMs like GPT-4 openai are queried in a zero-shot manner to decide the action sequentially zhou2023navgpt ; long2023discuss ; chen2024mapgpt ; chen2024affordances . For example, NavGPT zhou2023navgpt transforms visual observations into textual formats and feed them to LLMs for generating action predictions. The second branch finetunes open-source LLMs with in-domain VLN datasets, which alleviates the LLM’s query cost as well as mitigates the gap between LLM’s training corpus and VLN tasks zheng2024towards ; lin2024navcot ; zhang2024navid ; han2024roomtour3d . For example, Navid zhang2024navid constructs a video-based navigational vision-language model and train it using navigation samples collected from continuous R2R datasets. However, most of them map navigation inputs to action decisions without the reasoning output. In contrast, we train the open-source LLM to generate self-improving embodied reasoning explicitly to improve action decision accuracy, which enhances the decision interpretability as well as mitigates overfitting to training reasoning labels.
Embodied Chain-of-thoughts Training. Chain-of-thoughts (CoT) reasoning wei2022chain has been a widely utilized technique in Large Language Models (LLMs) and Vision-Language Models (VLMs). By generating the intermediate reasoning steps rather than directly predicting the answer, CoT can promote the answer accuracy for various tasks such as mathematical reasoning, commonsense reasoning, code generation, etc openai ; OpenAI_2023 ; liu2024deepseek . Inspired by this, some recent works have trained LLMs/VLMs to generate reasonable CoT for improving the action decision accuracy in embodied tasks zawalski2024robotic ; lin2024navcot ; mu2023embodiedgpt ; zhao2025cot ; liu2025spatialcot ; zhou2024navgpt . ECoT zawalski2024robotic trains vision-language-action models to generate embodied reasoning including object bounding boxes and end effector positions to encourage better adaptation for robotic manipulation tasks. NavGPT-2 zhou2024navgpt resorts to the GPT-4V model to collect free-form step-wise CoT reasoning data to improve the navigational reasoning ability of the VLN-specialized models Chen2022ThinkGA . CoT-VLA zhao2025cot incorporates explicit visual chain-of-thought reasoning into vision-language-action models by predicting future image frames before generating the action sequence. In our work, we propose a novel self-improving embodied CoT training paradigm for boosting the VLN, which effectively improves the reasoning ability and decision accuracy of the navigation model in various unseen scenarios. Moreover, we construct formalized CoT supervision, which can reduce redundant reasoning information and simultaneously promote the model’s inference speed.
Figure 2: Overview of our EvolveNav. EvolveNav introduces a two-stage training framework for fulfilling self-improving embodied reasoning. In Stage 1 Formalized CoT Supervised Fine-Tuning, the navigation agent is trained using pre-constructed formalized CoT labels to generate navigational reasoning by predicting the landmark needed to locate with the corresponding direction. In Stage 2 Self-Reflective Post-Training, the agent’s own reasoning outputs are introduced as the self-enriched CoT labels to enhance the supervision diversity. A self-reflective auxiliary task is also designed to guide the navigation agent to discriminate between correct and wrong reasoning outputs.
3 Method
In this section, we first introduce the problem definition of the VLN task (Sec. 3.1). Then, we present the model architecture of the LLM-based navigation agent in our EvolveNav (Sec. 3.2). Finally, we describe our proposed self-improving embodied reasoning framework in detail (Sec. 3.3). The overview of our EvolveNav is presented in Figure 2. Specifically, our EvolveNav consists of two stages: 1) Formalized CoT Supervised Fine-Tuning (Sec. 3.3.1), where the LLM-based VLN model is initially trained using formalized CoT labels with supervised fine-tuning to activate the model’s navigational reasoning ability and simultaneously promote the reasoning speed; 2) Self-Reflective Post-Training (Sec. 3.3.2), where the model is further trained with its own reasoning outputs as self-enriched CoT labels to increase supervision diversity, accompanied by a self-reflective auxiliary task to encourage better learning of accurate navigational reasoning patterns by discriminating from incorrect ones.
3.1 Problem Setup
In the VLN task, an agent is given a navigation instruction $I$ in the form of a declarative sentence or a dialogue and is required to navigate from a start position to the target position. At timestep $t$, the agent receives a panoramic observation $O_{t}$ containing $K$ single-view observations $O_{t,k}$, i.e., $O_{t}=\{O_{t,k}\}_{k=1}^{K}$. There are $N$ navigable views among $K$ views. The navigable views and the stop action form the action space, from which the agent chooses one as the action prediction $a_{t}$. Actions before step $t$ are treated as the navigation history.
3.2 Model Architecture
We build the LLM-based navigation agent that can simultaneously produce the navigational chain-of-thought reasoning and action prediction, modifying from a recent LLM-based VLN work, NaviLLM zheng2024towards . Concretely, the navigation agent consists of a scene encoder $F_{v}$, an LLM backbone $F_{\mathrm{LLM}}$, and an action prediction head $F_{\mathrm{action}}$. At each timestep $t$, the agent receives the navigation instruction $I$, panoramic observation $O_{t}$, and navigation history features $H_{t}=\{h_{0},...,h_{t-1}\}$. The scene encoder transforms $N$ navigable panoramic views $\{O_{t,n}\}_{n=1}^{N}$ into visual representations $\{V_{t,n}\}_{n=1}^{N}$:
$\{V_{t,n}\}_{n=1}^{N}$ $=F_{v}(\{O_{t,n}\}_{n=1}^{N}).$ (1)
The navigation prompt $P$ is then constructed by integrating the tokenized instruction, the visual representations $\{V_{t,n}\}_{n=1}^{N}$, and navigation history features $H_{t}$. As shown in Figure 2, special tokens <hist> and <cand> are introduced as placeholder tokens, where we insert the features $H_{t}$ and $\{V_{t,n}\}_{n=1}^{N}$, respectively. In contrast to NaviLLM zheng2024towards that directly maps the navigational inputs to action decision, in our EvolveNav, we construct the following output hint in the prompt $P$ to guide the navigation agent to generate both the action decision and explicit navigational reasoning: “-Action Decision: <cls>. -Navigational Reasoning: ”. The <cls> token is also a special token for facilitating subsequent action predictions. The navigation prompt $P$ is fed into the LLM backbone $F_{\mathrm{LLM}}$ to obtain the feature ${f}^{cls}_{t}$ of the <cls> token and the chain-of-thought (CoT) reasoning $\mathrm{CoT}$:
${f}^{cls}_{t},\mathrm{CoT}$ $=F_{\mathrm{LLM}}(P).$ (2)
Under the guidance of CoT reasoning, the ${f}^{cls}_{t}$ is sent to the action prediction head $F_{\mathrm{action}}$ for generating the action prediction $a_{t}$:
${a}_{t}$ $=F_{\mathrm{action}}({f}^{cls}_{t}).$ (3)
3.3 Self-Improving Embodied Reasoning
3.3.1 Stage 1: Formalized CoT Supervised Fine-Tuning
Formalized CoT Labels Collection. When facing a given human instruction, the navigation agent usually needs to sequentially reason about the direction or the landmark it should move to in its current visual observation to reach the target position. Therefore, in our EvolveNav, we train the LLM-based VLN model to generate the chain-of-thought (CoT) reasoning about the landmark with the corresponding direction at different navigation timesteps, like the following format: “I should go to an observation with [landmark] to the [direction] of me”.
To encourage the navigation agent to choose the ground-truth action $a^{*}_{t}$ (paired observation is denoted as $O_{t}^{*}$) at different timesteps $t$ through the guidance of CoT reasoning during training, we obtain the corresponding landmarks $L$ and direction $D$ of $O_{t}^{*}$ to construct the formalized CoT labels, which is described as follows. Denote the observation $O_{t}^{*}$ as $O_{t}^{*}=\{B_{t},A_{t}=\{\psi_{t},\theta_{t}\}\}$, where $B_{t}$ is the RGB image of $O_{t}^{*}$, $A_{t}$ represent the direction information containing heading $\psi_{t}$ and elevation $\theta_{t}$. For the image $B_{t}$, we first employ a powerful image captioning model huang2023tag2text $F_{\mathrm{cap}}$ to obtain object and scene context $C_{t}$:
$C_{t}=F_{\mathrm{cap}}(B_{t}).$ (4)
Then, we leverage the NLP tool Spacy spacy to extract the landmarks list $L$ from $C_{t}$. In contrast to directly using object recognition models which may detect multiple redundant objects, extracting landmarks from the image captions can better retain salient landmarks. As a result, the generated CoT reasoning of the navigation agent can effectively help it locate important landmarks mentioned in the human instruction, since humans also tend to focus on salient landmarks when giving navigation instructions. We follow lin2024navcot to map the direction information $A_{t}$ of the observation $O_{t}^{*}$ to textual represented direction $D$. With the landmarks list $L$ and direction $D$, we construct CoT labels $\mathrm{\mathrm{CoT}^{*}}$ by filling the following label template: I should go to an observation with [$L$] to the [$D$] of me. Through using our formalized CoT labels for training, the generation of redundant reasoning information can be effectively alleviated to promote the reasoning speed.
Supervised Fine-Tuning with Formalized CoT Labels. To activate the potential navigational reasoning ability of the LLM agent to adapt to the VLN task, we introduce the supervised-finetuning (SFT) paradigm in Stage 1 for conducting CoT training with our pre-constructed formalized CoT labels. Denote the navigation data sample at each timestep $t$ as $(P,\mathrm{CoT^{*}})$ (we omit the subscript $t$ for simplicity), where $P$ and $\mathrm{CoT^{*}}$ are the navigation prompt and CoT label, respectively. The training objective $\mathcal{L}_{\mathrm{SFT}}$ maximizes the likelihood of generating $\mathrm{CoT}^{*}$ given $P$ auto-regressively:
$\mathcal{L}_{\mathrm{SFT}}=-\mathbb{E}_{(P,\mathrm{CoT}^{*})\sim\mathcal{D}}\sum_{s=1}^{S}\mathrm{log}F_{LLM}(\mathrm{CoT}^{*}_{s}|P,\mathrm{CoT}^{*}_{<s}),$ (5)
where $\mathcal{D}$ represents the navigation dataset. Denote the navigation action prediction training objective as $\mathcal{L}_{\mathrm{action}}$, the total training objective $\mathcal{L}_{\mathrm{Stage1}}$ of Stage 1 is calculated as follows:
$\mathcal{L}_{\mathrm{Stage1}}=\mathcal{L}_{\mathrm{action}}+\lambda\mathcal{L}_{\mathrm{SFT}},$ (6)
where $\lambda$ represents the loss balance factor. We follow zheng2024towards to calculate the action prediction training objective $\mathcal{L}_{\mathrm{action}}$.
3.3.2 Stage 2: Self-Reflective Post-Training
Although CoT supervised fine-tuning can explicitly guide the navigation agent to produce navigational reasoning for assisting the action decision, due to the uncertainty and complexity of the navigation task, using fixed labels may lead to overfitting to training CoT label distributions and therefore harm the generalization to unseen scenarios. Moreover, the inherent noise in the image captioning model huang2023tag2text for landmark detection may also limit the accuracy of formalized CoT labels collected in Stage 1. Therefore, after the Stage 1 training of Formalized CoT Supervised Fine-Tuning, we introduce Self-Reflective Post-Training to further encourage the navigation agent to learn correct reasoning patterns for improving generalization.
Self-Enriched CoT Label Enhancement. To mitigate the overfitting to fixed CoT labels during training, we utilize the model’s self-generated reasoning outputs as self-enriched CoT labels under the guidance of the model’s action decision. At timestep $t$, denote the model’s reasoning output as $R_{t}$, the original formalized CoT label as $\mathrm{CoT}_{t}^{*}$, the model’s action decision as $a_{t}$, and the ground-truth action as $a_{t}^{*}$. When the action decision $a_{t}$ generated by the navigation agent matches the ground-truth action $a_{t}^{*}$, we choose the agent’s own reasoning output $R_{t}$ as the new CoT label. Such self-enriched CoT labels can effectively enhance the supervision diversity in a decision-oriented manner. Concretely, we obtain the updated CoT label $\tilde{\mathrm{CoT}}_{t}^{*}$ at timestep $t$ through the following rules:
$\tilde{\mathrm{CoT}}_{t}^{*}=\begin{cases}R_{t},\quad\quad\text{if}\quad a_{t}=a_{t}^{*}\\ \mathrm{CoT}_{t}^{*},\quad\text{otherwise}\end{cases}$ (7)
Self-Reflective Auxiliary Task. To further make the model aware of correct and wrong reasoning, which can help the model better learn correct reasoning patterns, we additionally introduce a self-reflective auxiliary task, where we ask the model to discriminate which reasoning output from the given reasoning is right. Specifically, we collect positive and negative reasoning samples $R^{+}$ and $R^{-}$ during training for conducting the self-reflective auxiliary task. We utilize the above mentioned CoT label $\tilde{\mathrm{CoT}}_{t}^{*}$ as the positive reasoning sample $R^{+}$. To obtain the negative reasoning sample $R^{-}$, we randomly select the candidate observation (action) $O_{t,j}$ ($1<j<N$, $N$ is the number of navigable views) different from the ground-truth one. Then we extract the landmark and direction for $O_{t,j}$ to fill in the CoT label template (see Sec. 3.3.1). As shown in Figure 2, we construct the self-reflective task prompt $P_{\mathrm{sr}}$ as “Choose the correct one from the given two navigational reasoning outputs. Output 1: [$R^{1}$]. Output 2: [$R^{2}$]. Selection: ”, where we randomly insert the positive reasoning sample $R^{+}$ and negative reasoning sample $R^{-}$ to the positions of $R^{1}$ and $R^{2}$. We collect the ground-truth output $R_{\mathrm{sr}}^{*}$ with the form like “Output 2.”. We also utilize the auto-regressive training objective like Eq. 5 to calculate the loss $\mathcal{L}_{\mathrm{sr}}$ based on ($P_{\mathrm{sr}}$, $R_{\mathrm{sr}}^{*}$) pairs for the self-reflective auxiliary task.
With the self-reflective loss $\mathcal{L}_{\mathrm{sr}}$, the total training objective $\mathcal{L}_{\mathrm{Stage2}}$ for Stage 2 is obtained by:
$\mathcal{L}_{\mathrm{Stage2}}=\mathcal{L}_{\mathrm{action}}+\lambda_{1}\mathcal{L}_{\mathrm{SFT}}+\lambda_{2}\mathcal{L}_{\mathrm{sr}},$ (8)
where both $\lambda_{1}$ and $\lambda_{2}$ are the loss coefficients. Through the self-enriched CoT label enhancement and self-reflective auxiliary task, the navigation agent learns to generate correct embodied reasoning in a self-refining manner to enable adaptable reasoning in different scenarios.
4 Experiment
4.1 Experimental Setup
Datasets. We test our EvolveNav on two popular VLN benchmarks, i.e., R2R anderson2018vision and CVDN thomason2019vision . R2R is built on 90 real-world indoor simulation environments containing 7,189 trajectories, each corresponding to three fine-grained instructions. CVDN contains 2,050 human-human navigation dialogs and over 7k trajectories in 83 MatterPort houses.
Evaluation Metrics. We utilize the following standard metrics for evaluation: 1) Trajectory Length (TL): the average length of the agent’s navigated path, 2) Navigation Error (NE): the average distance between the agent’s destination and the goal viewpoint, 3) Success Rate (SR): the ratio of success, where the agent stops within three meters of the target point, 4) Success rate weighted by Path Length (SPLanderson2018vision : success rate normalized by the ratio between the length of the shortest path and the predicted path, 5) Oracle Success Rate (OSR): the ratio of containing a viewpoint along the path where the target position is visible, 6) Goal Progress (GP), the progress in meters towards the goal.
Table 1: Comparison with state-of-the-art methods on the R2R dataset. * denotes training with R2R in-domain data. IL means the imitation learning setting.
Method Val Unseen Test Unseen
[…]
Cross-Modal Backbone:
PREVALENT hao2020towards 10.19 4.71 58 53 10.51 5.30 54 51
HOP qiao2022hop 12.27 3.80 64 57 12.68 3.83 64 59
HAMT Chen2021HistoryAM 11.46 2.29 66 61 12.27 3.93 65 60
VLN-BERT hong2021vln 12.01 3.93 63 57 12.35 4.09 63 57
DUET Chen2022ThinkGA 13.94 3.31 72 60 14.73 3.65 69 59
Meta-Explore Hwang2023MetaExploreEH 13.09 3.22 72 62 14.25 3.57 71 61
VLN-SIG li2023improving - - 72 62 - - 72 60
VLN-PETL qiao2023vln 11.52 3.53 65 60 12.30 4.10 63 58
NavGPT2 zhou2024navgpt 13.25 3.18 71 60 - - - -
LLM-based Backbone:
NavGPT zhou2023navgpt 11.45 6.46 34 29 - - - -
DiscussNav long2023discuss 9.69 5.32 43 40 - - - -
MapGPT chen2024mapgpt - 5.63 34 29 - - - -
NavCoT lin2024navcot 9.95 6.26 40 37 - - - -
NaviLLM zheng2024towards 12.81 3.51 67 59 13.21 3.71 68 60
NaviLLM* (IL) 9.99 6.04 46.90 43.78 10.03 6.12 46 43
EvolveNav (IL) 9.79 5.52 51.15 48.27 9.94 5.92 47 45
NaviLLM* 13.43 3.27 70.11 60.25 13.68 3.37 70 61
EvolveNav 12.07 3.15 71.17 63.48 12.06 3.22 71 63
Implementation Details. We fine-tune the LLM with full-parameter and LoRA settings for Stage 1 and 2, respectively. To simplify the implementation, we do not introduce the pretraining phase like zheng2024towards for both our EvolveNav and the compared baseline (denoted as NaviLLM* in Table 1 &  2). The training for Stage 1 is conducted on 8 Nvidia A100 GPUs and the training for Stage 2 is performed on 4 Nvidia A100 GPUs. Empirically, we set the loss coefficients $\lambda$, $\lambda_{1}$, and $\lambda_{2}$ as 1, 1, and 0.2, respectively. The hyperparameters such as the learning rate, optimizer, and the batch size are kept the same as zheng2024towards .More implementation details are presented in the supplementary material.
4.2 Comparison with Existing Methods
Table 1 shows the comparison among different approaches on R2R. We compare EvolveNav with the baseline method NaviLLM zheng2024towards using the R2R in-domain training data (denoted as NaviLLM*) under two settings, i.e., imitation learning (IL) and Dagger training zheng2024towards ; Chen2022ThinkGA . The comparison shows the consistent improvement of our approach over NaviLLM* on both Val Unseen and Test Unseen under different training settings. For example, the performance gain in SPL of EvolveNav on Val Unseen is $\sim$4.5and $\sim$3.2under IL and dagger training settings compared to the baseline, respectively. Note that our EvolveNav achieves the state-of-the-art performance on R2R among the approaches with LLM-based backbone. These results demonstrate the effectiveness of our self-improving embodied reasoning framework in fulfilling generalizable navigational reasoning. We do not consider the comparison with models augmented by new environments (e.g., ScaleVLN wang2023scaling ) for fairness.
Table 2: Comparison with SoTA methods on CVDN. We utilize the Goal Progress (GP) (m) as the evaluation metric. * denotes training with CVDN in-domain data.
Method Val-Unseen Test
Cross-Modal Backbone:
Seq2Seq thomason2019vision 2.10 2.35
PREVALENT hao2020towards 3.15 2.44
HOP qiao2022hop 4.41 3.31
MT-RCM wang2020environment 4.36 -
MT-RCM+Env wang2020environment 4.65 3.91
HAMT Chen2021HistoryAM 5.13 5.58
VLN-SIG li2023improving 5.52 5.83
VLN-PETL qiao2023vln 5.69 6.13
LLM-based Backbone:
NaviLLM zheng2024towards 6.16 7.90
NaviLLM* 5.53 6.80
EvolveNav (ours) 6.21 7.07
In Table 2, we compare our EvolveNav with different VLN approaches on CVDN. As shown in Table 2, EvolveNav outperforms the baseline NaviLLM* zheng2024towards on both Val Unseen and Test set on CVDN, demonstrating good generalization ability of our proposed self-improving embodied reasoning paradigm to different kinds of VLN tasks. For example, EvolveNav brings $\sim$0.7 performance gain over NaviLLM* in GP on Val Unseen. Moreover, Table 2 shows the superiority of EvolveNav over NaviLLM zheng2024towards , which conducts multi-task training on multiple VLN benchmarks to learn cross-task navigation knowledge, demonstrating the importance of explicit navigational reasoning learning compared to direct input-output mapping paradigms. Such results also reveal the effectiveness of our constructed CoT labels, which provide landmark and direction information for enabling the agent to learn useful navigational reasoning for decisions. The results on both R2R and CVDN benchmarks demonstrate the effectiveness of our EvolveNav in fulfilling generalizable navigational reasoning to improve navigation decision accuracy.
4.3 Ablation Study
Effect of different method components. Table 3 presents the ablation study on Val Unseen set on R2R, where we can find the effectiveness and reasonability of different method components in our EvolveNav. From Table 3, we can observe that through the Stage 1 training of formalized CoT supervised fine-tuning (SFT), the model’s navigational reasoning ability can be significantly enhanced to improve the navigation performance. The introduction of self-enriched CoT labels and the self-reflective auxiliary task can further bring performance gain in both SR and SPL metrics compared to pure CoT SFT. Our full model achieves the best results in SR and SPL among all variants, demonstrating the effectiveness of our two-phase self-improving embodied training framework.
Table 3: Ablation study of method components on Val Unseen set on R2R. We adopt the imitation learning (IL) setting for evaluation.
[…]
Baseline 46.90 54.63 43.78
   + Formalized CoT SFT 49.62 59.44 46.26
   + Self-Enriched CoT SFT 50.47 57.48 47.98
   + Self-reflective task 50.51 57.74 47.86
   + Full Model (ours) 51.15 59.18 48.27
Effect of constructed CoT labels. Table 4 compares the navigation performance under different kinds of CoT labels, where we can find the effectiveness of our introduced CoT labels by predicting landmark and direction information in a formalized way. To realize “Free-form CoT”, we introduce the free-form CoT labels collected in NavGPT-2 zhou2024navgpt to train the navigation agent. To realize “Direction & Landmark”, we obtain the best matched landmark in the instruction to each ground-truth observation through the CLIP model radford2021learning . The comparison between “Free-form CoT” and “Direction & Landmark (ours)” shows that our formalized CoT labels can effectively reduce redundant reasoning information to improve the navigational reasoning and decision accuracy. Moreover, the inference time to generate CoT at one timestep of “Free-form CoT” is $\sim$7.8s compared to $\sim$2.5s of “Direction & Landmark (ours)”, demonstrating that our method can significantly promote the reasoning speed ($\sim$×$\times$3 improvement), which is crucial for sequential decision making task like navigation. By comparing “Only Direction” and “Only Landmark” with “Direction & Landmark (ours)”, we can find that both landmark and direction information are important to guide navigation decisions, revealing the reasonability of our constructed CoT labels. The superiority of “Direction & Landmark (ours)” over “Direction & Landmark” demonstrates that our introduced CoT labels, which contain diverse landmarks, can potentially encourage the navigation agent to learn cross-modal alignment knowledge to accurately align the visual observation to the navigation instruction.
Table 4: Ablation study of CoT on Val Unseen set on R2R. denotes using the landmark mentioned in the instruction.
[…]
Baseline 70.11 79.00 60.25
Free-form CoT 69.39 80.44 60.95
Only Direction 67.05 76.49 57.32
Only Landmark 68.32 79.85 60.53
Direction & Landmark 69.77 79.25 60.77
Direction & Landmark (ours) 71.26 80.23 62.05
4.4 Visualization
Fig. 3 gives the action decision visualization comparison between NaviLLM zheng2024towards and our EvolveNav, where we can find that EvolveNav generates reasonable navigational reasoning about landmarks and directions to guide correct action decision making. For example, when $t=0$, from the observations, EvolveNav infers that an observation with door and hallway represents the exit from the bedroom while NaviLLM mistakenly selects an action that remains in the bedroom. Another example is in the hallway ($t=4$ for NaviLLM and $t=1$ for EvolveNav), EvolveNav generates reasoning consistent with the decision of entering the correct doorway on the left. However, NaviLLM chooses the wrong doorway to go back. These results highlight the effectiveness of our approach in improving navigational reasoning for accurate instruction understanding and action prediction.
Figure 3: Action decision visualization of NaviLLM zheng2024towards and our EvolveNav. We only extract two steps and display local candidate space for simplicity. Observations selected by EvolveNav (also are the ground-truth actions) and NaviLLM are annotated by green boxes and red boxes, respectively.
5 Conclusion
In this work, we propose EvolveNav, a novel self-improving embodied reasoning framework to fulfill generalizable and adaptable reasoning for enhancing LLM-based vision-and-language navigation. Through introducing the formalized CoT supervised fine-tuning and self-reflective post-training in the proposed framework, the agent’s navigational reasoning ability can be effectively enhanced while mitigating the overfitting to the training reasoning label distributions simultaneously to improve generalization. Experimental results on two popular benchmarks R2R and CVDN reveal the promising capability of our method in boosting the reasoning ability and decision accuracy for LLM-based navigation agents. We believe that our EvolveNav can provide meaningful references for designing self-improving embodied reasoning paradigms to benefit future LLM-assisted Embodied AI research.
Supplementary Material
Appendix A Overview
In these supplementary materials, we first analyze the broader impacts and limitations of our EvolveNav in Sec. B and Sec. C, respectively. Then, we describe more implementation details in Sec. D. Finally, we present more experimental results, including the landmark extraction visualization, self-enriched CoT label visualization, and loss & performance curves during Self-Reflective Post-Training, in Sec. E.
Appendix B Broader Impact
The scalable deployment of LLM-based navigation robots holds significant potential to enhance operational efficiency in public spaces, reduce human labor costs, and benefit diverse populations including the elderly and individuals with disabilities. Our proposed self-improving embodied reasoning framework can help improve both navigation success rate and decision interpretability of LLM-based navigation agents, therefore contributing to scalable application of LLM-based navigation robots which have high robustness and reliability. However, the inherent hallucination of LLMs may cause incorrect decisions in complex environments, leading to potential risks such as collisions. More efforts devoted to hallucination suppression of LLMs and safety promotion of navigation robots are encouraged for future research in the community.
Appendix C Limitations
(1) In the Stage 1 training of Formalized CoT Supervised Fine-tuning in our EvolveNav, we conduct full-parameter training for the open-source LLM, vicuna-7B [1], like the baseline method [66]. This process costs $\sim$1.5 day training on 8 NVIDIA A100 GPUs with 80G memory, which is relative resource-intensive compared to conventional VLN approaches [10, 9, 17]. In future work, we plan to introduce parameter-efficient training strategies or resort to more lightweight LLMs to reduce the computational resource consumption.
(2) During the formalized CoT label collection for supervised fine-tuning, we employ the open-vocabulary image captioning model [24] to obtain the caption for visual observations, followed by the NLP tool Spacy [20] to extract the landmark from the caption. This process is performed in an offline manner, which might introduce noise for the CoT label construction due to the possible inaccuracy of the image captioning model and the NLP tool. Future work includes utilizing more powerful open-vocabulary models or conducting human checking to promote the accuracy of the CoT labels.
Appendix D Implementation Details
We adopt Vicuna 7B [1] as our large language model (LLM) backbone, following [66]. For R2R [5], we also introduce the auxiliary tasks of trajectory summarization and embodied question answering in [66] during training. The maximum number of landmark extracted for constructing formalized CoT labels is set as 5 to retain useful information and simultaneously promote the reasoning speed. To mitigate the possible landmark missing during landmark detection due to the view deviation for a single candidate observation, we fuse the landmark information of candidate observations with the same direction to construct the CoT labels. We remove some uninformative landmarks such as “floor” and “ceiling” empirically during the landmark extraction. During Stage 1 training, we introduce the CoT supervised finetuning loss $\mathcal{L}_{\mathrm{SFT}}$ under a probability of 0.5 to mitigate the overfitting to the pre-constructed CoT labels. The maximum numbers of training steps for Stage 1 and 2 are set as 60000 and 9000 steps, respectively. Training for Stage 1 with full parameter lasts for $\sim$1.5 days with $\sim$73G GPU memory, and training for Stage 2 with LoRA lasts for $\sim$1 day with $\sim$30G GPU memory. The inference is conducted on 4 V100 GPUs with the evaluation batch size of 2. During inference, we set the LLM temperature to 0 and opt for a greedy strategy to generate action decisions.
Appendix E More Experimental Results
Landmark extraction visualization. We compare different methods for constructing the formalized CoT labels, to verify the reasonability of our landmark extraction strategy by combining image captioning model [24] with the NLP tool Spacy [20]. Fig. 4 presents the landmark extraction visualization comparison of three image captioning models, Tag2Text [24], BLIP-v2 [29], LLaVA 1.6 vicuna 13b [35], and an open-vocabulary object recognition model RAM [64]. Concretely, we provide four candidates in a navigational step for these methods and use Spacy [20] to extract the landmarks in the output caption (except for RAM of directly obtaining tagging). From Fig. 4, we can observe that Tag2Text can capture more informative landmarks while having less redundancy and illusion. For example, for “Cand 2”, Tag2Text correctly detects a kitchen and living room, a dining table, and chair, while BLIP-v2 only detects a large dining table. Although LLaVA 1.6 13b generates abundant captions, it brings noisy landmarks like the outdoor after noun phrases extraction. RAM also generates meaningless and non-existent landmarks, like floor and armchair. These results show that our combination of Tag2text model [24] and NLP tool for landmark extraction can effectively retain informative landmarks while reducing redundancy and illusion for constructing CoT labels, which can enable the agent to better learn cross-modal alignment between observations and instructions.
Figure 4: Landmark extraction visualization of different methods. We use green, red, and blue colors to distinguish informative, false, and uninformative landmarks, respectively.
Self-Enriched CoT label visualization. In Fig. 5, we present some visualization comparison examples of original CoT labels and self-enriched CoT labels. From Fig. 5, we can observe that our self-enriched CoT label enhancement strategy effectively increases the supervision diversity. For example, in Fig. 5(a) and (b), the self-enriched CoT labels from the model’s own reasoning outputs capture the landmark painting and chair which are not contained in the original CoT label, respectively. From Fig. 5(c) and (d), we can find that the LLM-based navigation agent can also recognize the attribute of the landmark, e.g., it recognizes wood paneling and arched window, which indicate the material and shape of the landmark, respectively. Such CoT labels can help navigation agent learn to follow more fine-grained instructions. Benefiting from our self-enriched CoT label, the LLM-based navigation agent can reduce the overfitting to the original CoT label distributions and learn more diverse cross-modal alignment knowledge, and therefore promote the generalization to unseen scenarios.
Figure 5: Self-enriched chain-of-thought (CoT) label visualization. Newly introduced landmarks in the self-enriched CoT label are highlighted in red fonts. GT action denotes the ground-truth action (observation). We omit the direction information in the CoT label.
(a) Loss variation
(b) Navigation Error (NE) variation
Figure 6: Loss and performance variation during Stage 2: Self-Reflective Post-Training. Low navigation error (NE) value indicates better results.
Loss & performance variation during Stage 2 training. Fig. 6 shows the loss and performance curves during Self-Reflective Post-Training (Stage 2). In Fig. 6, we can find that our self-reflective loss $\mathcal{L}_{\mathrm{sr}}$, the total training loss $\mathcal{L}_{\mathrm{Stage2}}$, and the navigation error (NE) have similar variation trends. Especially, both two loss curves and the NE curves achieve the lowest value around epoch 9. These results show the effect of our self-reflective auxiliary task during Self-Reflective Post-Training in improving the navigational reasoning and decision accuracy of the LLM-based navigation agent.

### v3

@@
-EvolveNav: Self-Improving Embodied Reasoning for LLM-Based Vision-Language Navigation
+EvolveNav: Empowering LLM-Based Vision-Language Navigation via Self-Improving Embodied Reasoning
 Abstract
-Building Vision-Language Navigation (VLN) agents which can navigate following natural language instructions is a long-standing goal in human-robot interaction applications.
-Recent studies have revealed the potential of training open-source Large Language Models (LLMs) to unleash LLMs’ reasoning ability for improving navigation, and simultaneously mitigate the domain gap between LLMs’ training corpus and the VLN task.
-However, these approaches primarily adopt direct input-output mapping paradigms, causing the mapping learning difficult and the navigational decisions unexplainable.
+Recent studies have revealed the potential of training open-source Large Language Models (LLMs) to unleash LLMs’ reasoning ability for enhancing vision-language navigation (VLN) performance, and simultaneously mitigate the domain gap between LLMs’ training corpus and the VLN task.
+However, these approaches predominantly adopt straightforward input-output mapping paradigms, causing the mapping learning difficult and the navigational decisions unexplainable.
 Chain-of-Thought (CoT) training is a promising way to improve both navigational decision accuracy and interpretability, while the complexity of the navigation task makes the perfect CoT labels unavailable and may lead to overfitting through pure CoT supervised fine-tuning.
-In this paper, we propose a novel sElf-improving embodied reasoning framework for boosting LLM-based vision-language Navigation, dubbed EvolveNav.
-Our EvolveNav consists of two stages: (1) Formalized CoT Supervised Fine-Tuning, where we train the model with formalized CoT labels to both activate the model’s navigational reasoning capabilities and increase the reasoning speed; (2) Self-Reflective Post-Training, where the model is iteratively trained with its own reasoning outputs as self-enriched CoT labels to enhance the supervision diversity.
-A self-reflective auxiliary task is also introduced to encourage learning correct reasoning patterns by contrasting with wrong ones.
-Experimental results on the popular VLN benchmarks demonstrate the superiority of EvolveNav over previous LLM-based VLN approaches.
-Code is available at: https://github.com/expectorlin/EvolveNav.
+To address these issues, we propose EvolveNav, a novel sElf-improving embodied reasoning paradigm that realizes adaptable and generalizable navigational reasoning for boosting LLM-based vision-language Navigation.
+Specifically, EvolveNav involves a two-stage training process: (1) Formalized CoT Supervised Fine-Tuning, where we train the model with curated formalized CoT labels to first activate the model’s navigational reasoning capabilities, and simultaneously increase the reasoning speed; (2) Self-Reflective Post-Training, where the model is iteratively trained with its own reasoning outputs as self-enriched CoT labels to enhance the supervision diversity.
+A self-reflective auxiliary task is also designed to encourage the model to learn correct reasoning patterns by contrasting with wrong ones.
+Experimental results under both task-specific and cross-task training paradigms demonstrate the consistent superiority of EvolveNav over previous LLM-based VLN approaches on various popular benchmarks, including R2R, REVERIE, CVDN, and SOON.
+EvolveNav open avenues for exploring effective self-improving reasoning paradigms, enabling building agents capable of self-evolving for promoting LLM-based embodied AI research.
+publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
 1 Introduction
@@
 In VLN tasks, an embodied agent needs to follow natural language instructions to navigate through complex visual environments to reach the target position.
-Early works improve VLN performance by designing dedicated model architectures wang2019reinforced ; ma2019self ; deng2020evolving ; qi2020Object , introducing powerful learning paradigms zhu2020vision ; li2019robust ; huang2019transferable , and developing useful data augmentation techniques tan2019learning ; fried2018speaker ; liu2021vision ; Fu2019CounterfactualVN .
-Subsequently, pretraining-based VLN approaches have been widely proposed to improve the cross-modal alignment ability and decision accuracy of navigation agents hong2021vln ; Chen2021HistoryAM ; Qi2021TheRT ; Chen2022ThinkGA ; Qiao2022HOPHA .
+Early works improve VLN performance by designing dedicated model architectures [1, 2, 3, 4], introducing powerful learning paradigms [5, 6, 7], and developing useful data augmentation techniques [8, 9, 10, 11].
+Subsequently, pretraining-based VLN approaches have been widely proposed to improve the cross-modal alignment ability and decision accuracy of navigation agents [12, 13, 14, 15, 16].
 Nevertheless, constrained by the limited scale of pretraining and VLN in-domain data, these approaches cannot learn navigational reasoning knowledge sufficiently, and therefore still struggle to handle various unseen navigation scenarios.
-With the rapid progress of large language models (LLMs) brown2020language ; touvron2023llama ; touvron2023llama2 , emerging works have introduced LLMs to address embodied tasks considering LLMs’ rich real-world common sense and powerful reasoning ability Ahn2022DoAI ; Huang2022InnerME ; driess2023palme .
-Some recent works have attempted to build LLM-based VLN models in a zero-shot or trainable manner  chen2024mapgpt ; zhou2023navgpt ; long2023discuss ; hao2020towards ; zhang2024navid .
-The zero-shot approaches, such as NavGPT zhou2023navgpt and MapGPT chen2024mapgpt , resort to closed-source LLMs OpenAI_2023 to generate navigational reasoning and decision for different navigation timesteps.
-To alleviate the high cost of frequently querying closed-source LLMs for sequential decision making, the trainable methods collect in-domain data to train open-source LLMs vicuna to build the navigation agent.
+With the rapid progress of large language models (LLMs) [17, 18, 19], emerging works have introduced LLMs to address embodied tasks by resorting to LLMs’ rich real-world common sense and powerful reasoning ability [20, 21, 22].
+Some recent works have attempted to build LLM-based VLN models in a zero-shot or trainable manner  [23, 24, 25, 26, 27].
+The zero-shot approaches, such as NavGPT [24] and MapGPT [23], resort to closed-source LLMs [28] to generate navigational reasoning and decision for different navigation timesteps.
+To alleviate the high cost of frequently querying closed-source LLMs for sequential decision making, the trainable methods collect in-domain data to train open-source LLMs [29] to build the navigation agent.
 However, they typically map directly from navigational inputs to decisions without explicit intermediate reasoning steps, leading to decision uninterpretability and may also limit performance (see Figure 1(a)).
@@
 R+ and R- represent positive and negative reasoning samples, respectively.
-Recent studies have revealed the effectiveness of embodied chain-of-thoughts (CoT) training in enhancing both the decision accuracy and interpretability for embodied tasks zawalski2024robotic ; zhao2025cot ; lin2024navcot .
-Most of these approaches employ the supervised fine-tuning paradigm for embodied CoT training.
-However, introducing CoT supervised fine-tuning for training VLN models is challenging.
-Firstly, due to the complexity and uncertainty of the navigation task, there can be multiple cues for deciding the correct navigation action, i.e., there can be multiple “correct” CoT supervisions to guide navigation for a specific timestep.
-This leads to a hard collection process of perfect navigation CoT labels.
-Secondly, pure CoT supervised fine-tuning using fixed CoT labels may cause overfitting to certain reasoning patterns and harm the generalization to diverse unseen scenarios.
-In this paper, we propose a sElf-improving embodied reasoning paradigm for enhancing LLM-based vision-language Navigation, called EvolveNav, to fulfill generalizable and adaptable navigational reasoning under different scenes.
-Our EvolveNav comprises two stages: 1) Formalized CoT Supervised Fine-Tuning and 2) Self-Reflective Post-Training.
-The Stage 1 training of Formalized CoT Supervised Fine-Tuning aims to first activate the model’s potential reasoning capabilities, where we ask the model to produce explicit chain-of-thought navigational reasoning by predicting the landmarks needed to locate with the corresponding direction for deciding the navigation actions.
-To alleviate redundant reasoning and increase the inference speed, we conduct the CoT supervised fine-tuning using formalized CoT labels, which are collected by filling the landmark and direction information into concise label templates.
-To mitigate the overfitting to pre-constructed CoT labels and enable self-improving reasoning for enhancing generalization, we further conduct Self-Reflective Post-Training for Stage 2 training.
-Concretely, we design a self-enriched CoT label enhancement strategy, where we train the model with its iteratively produced correct reasoning outputs to enrich the CoT supervision.
-We also introduce a self-reflective auxiliary task, where the model needs to discriminate between positive and negative navigational reasoning to learn correct reasoning patterns better.
+Recent studies have revealed the effectiveness of chain-of-thoughts (CoT) training in enhancing both the decision accuracy and interpretability for embodied tasks [30, 31, 32].
+Most of these approaches employ the supervised fine-tuning (SFT) paradigm for conducting embodied CoT training.
+However, introducing CoT supervised fine-tuning for training VLN models is highly challenging due to the following two reasons.
+Firstly, due to the complexity and uncertainty of the navigation task, there can be multiple cues for deciding the correct navigation action, i.e., there may be no single “correct” CoT label to guide navigation for a specific timestep.
+This leads to a hard collection process of perfect navigation CoT supervision.
+Secondly, pure CoT supervised fine-tuning using fixed CoT labels may cause overfitting to certain reasoning patterns and thus harm the generalization to diverse unseen scenarios.
+In this paper, we propose a novel sElf-improving embodied reasoning paradigm for enhancing LLM-based vision-language Navigation, called EvolveNav, to fulfill generalizable and adaptable navigational reasoning under various tasks and scenarios.
+EvolveNav comprises two training phases: 1) Formalized CoT Supervised Fine-Tuning and 2) Self-Reflective Post-Training.
+The Stage 1 training of Formalized CoT Supervised Fine-Tuning aims to first activate the model’s potential reasoning capabilities, where we ask the model to produce explicit chain-of-thought navigational reasoning dynamically by predicting the landmarks needed to locate with the corresponding direction for deciding the navigation actions.
+To alleviate generating redundant reasoning and increase the inference speed, we conduct the CoT supervised fine-tuning using curated formalized CoT labels, which are collected by filling the landmark and direction information into concise label templates.
+Then, we conduct Self-Reflective Post-Training for Stage 2 training, aiming to mitigate the overfitting to pre-constructed CoT labels and enable self-improving reasoning for enhancing generalization.
+Specifically, we design a self-enriched CoT label enhancement scheme, where we train the model with its iteratively produced correct reasoning outputs to diversify the CoT supervision.
+We also construct a self-reflective auxiliary task, where the model needs to discriminate between positive and negative navigational reasoning to learn correct reasoning patterns.
 As shown in Figure 1, in contrast to CoT supervised fine-tuning using fixed CoT labels (i.e., Figure 1(b) Formalized CoT Learning and (c) Free-form CoT Learning), our EvolveNav (Figure 1(d)) can generate embodied CoT in a self-refining manner during training to mitigate the overfitting.
 Additionally, through training with formalized CoT labels, our EvolveNav can significantly reduce uninformative navigational reasoning to promote the reasoning speed compared with using free-form CoT labels for training (as in Figure 1 (c)).
-We conduct experiments on two public VLN benchmarks, R2R anderson2018vision and CVDN thomason2019vision .
-Experimental results show that EvolveNav substantially outperforms previous LLM-based VLN approaches on unseen scenarios, demonstrating the effectiveness of our self-improving embodied reasoning paradigm in promoting navigation decision accuracy and generalization.
+We conduct substantial experiments under both task-specific and cross-task training paradigms on multiple public VLN benchmarks, including R2R [33], CVDN [34], REVERIE [35], and SOON [36].
+Experimental results show that EvolveNav significantly outperforms previous LLM-based VLN approaches on various benchmarks, demonstrating the effectiveness of our self-improving embodied reasoning paradigm in promoting navigation decision accuracy and generalization.
 We carefully conduct ablation experiments to explore how to design streamlined CoTs that can provide interpretability while boosting navigation performance.
@@
 To summarize, the main contributions of this paper are:
-  • We propose a novel self-improving embodied reasoning paradigm, EvolveNav, for enhancing the LLM-based vision-and-language navigation, which fulfills generalizable and adaptable navigational reasoning under unseen scenes.
+  • We propose EvolveNav, a novel self-improving embodied reasoning paradigm for enhancing LLM-based vision-and-language navigation, which fulfills generalizable and adaptable navigational reasoning under various tasks and scenarios.
   • We construct formalized CoT labels for conducting supervised fine-tuning, which effectively activates the agent’s navigational reasoning ability and promotes the reasoning speed.
 We introduce a self-enriched CoT label enhancement strategy and a self-reflective auxiliary task to enable learning correct reasoning patterns in a self-refining manner to mitigate overfitting.
-  • Experimental results show the superiority of EvolveNav over other LLM-based approaches on two standard VLN benchmarks R2R and CVDN.
-Our EvolveNav can improve the generalization of both navigational reasoning and decision-making, providing meaningful insights for designing further embodied reasoning paradigms.
+  • Experimental results demonstrate the superiority of EvolveNav over previous LLM-based approaches on various VLN benchmarks.
+Our EvolveNav can improve the generalization of both navigational reasoning and decision-making, providing meaningful insights for designing advanced embodied reasoning paradigms.
 2 Related Work
-Vision-Language Navigation (VLN).
+2.1 Vision-Language Navigation (VLN)
 Vision-Language Navigation (VLN) has attracted intensive research interest in recent years.
-Various VLN benchmarks have been proposed to evaluate agents’ ability for navigational reasoning and instruction following anderson2018vision ; jain2019stay ; ku2020room ; thomason2019vision ; qi2020reverie ; zhu2021soon .
-Previous approaches employ non-pretraining-based wang2019reinforced ; ma2019self ; deng2020evolving ; zhu2020vision ; tan2019learning ; fried2018speaker or pretraining-based paradigms hong2021vln ; Chen2021HistoryAM ; Chen2022ThinkGA ; Qiao2022HOPHA ; Guhur2021AirbertIP ; an2022bevbert ; wang2023scaling for tackling the above VLN tasks.
+Various VLN benchmarks have been proposed to evaluate agents’ ability for navigational reasoning and instruction following [33, 37, 38, 34, 35, 36].
+Previous approaches employ non-pretraining-based [1, 2, 3, 5, 8, 9] or pretraining-based paradigms [12, 13, 15, 16, 39, 40, 41] for tackling the above VLN tasks.
 However, these approaches cannot generalize well to diverse unseen scenarios that require rich real-world commonsense, and the navigation decisions also lack explainability.
-Some recent works have introduced LLMs to assist the VLN task, by either eliciting the useful navigation knowledge stored in LLMs lin2024correctable ; zhou2024navgpt ; qiao2023march or employing the LLM as the navigation backbone for action decision zhou2023navgpt ; zheng2024towards ; zhang2024navid ; lin2024navcot ; chen2024mapgpt .
+Some recent works have introduced LLMs to assist the VLN task, by either eliciting the useful navigation knowledge stored in LLMs [42, 43, 44] or employing the LLM as the navigation backbone for action decision [24, 45, 27, 32, 23].
 Our work lies in the latter.
@@
 As a result, both the reasoning ability and decision interpretability of the navigation model can be significantly enhanced.
-LLMs as Embodied Agents.
-Recent research have revealed the giant potential of utilizing Large Language Models (LLMs) as embodied agents to complete the robotic navigation and manipulation tasks, benefiting from the outstanding ability of planning, reasoning, and reflection of LLMs song2023llmplanner ; Ahn2022DoAI ; Huang2022InnerME ; Yao2022ReActSR ; shahlm ; schumann-2023-velma ; wang2023voyager ; huang2023voxposer ; rajvanshi2024saynav .
-For example, LM-Nav shahlm introduces the LLM to parse the long navigation instruction into sequential landmarks for facilitating the navigational planning.
-Voxposer huang2023voxposer introduces LLMs for code writing and combines them with the Vision-Language models (VLMs) to compose 3D value maps for robotic manipulation.
+2.2 LLMs as Embodied Agents
+Recent research have revealed the giant potential of utilizing Large Language Models (LLMs) as embodied agents to complete the robotic navigation and manipulation tasks, benefiting from the outstanding ability of planning, reasoning, and reflection of LLMs [46, 20, 21, 47, 48, 49, 50, 51, 52].
+For example, LM-Nav [48] introduces the LLM to parse the long navigation instruction into sequential landmarks for facilitating the navigational planning.
+Voxposer [51] introduces LLMs for code writing and combines them with the Vision-Language models (VLMs) to compose 3D value maps for robotic manipulation.
 There are typically two branches of works where the LLMs act as embodied agents for tackling the VLN task.
-In the first branch, closed-source LLMs like GPT-4 openai are queried in a zero-shot manner to decide the action sequentially zhou2023navgpt ; long2023discuss ; chen2024mapgpt ; chen2024affordances .
-For example, NavGPT zhou2023navgpt transforms visual observations into textual formats and feed them to LLMs for generating action predictions.
-The second branch finetunes open-source LLMs with in-domain VLN datasets, which alleviates the LLM’s query cost as well as mitigates the gap between LLM’s training corpus and VLN tasks zheng2024towards ; lin2024navcot ; zhang2024navid ; han2024roomtour3d .
-For example, Navid zhang2024navid constructs a video-based navigational vision-language model and train it using navigation samples collected from continuous R2R datasets.
-However, most of them map navigation inputs to action decisions without the reasoning output.
+In the first branch, closed-source LLMs like GPT-4 [53] are queried in a zero-shot manner to decide the action sequentially [24, 25, 23, 54].
+For example, NavGPT [24] transforms visual observations into textual formats and feed them to LLMs for generating action predictions.
+The second branch finetunes open-source LLMs with in-domain VLN datasets, which alleviates the LLM’s query cost as well as mitigates the gap between LLM’s training corpus and VLN tasks [45, 32, 27, 55].
+For example, Navid [27] constructs a video-based navigational vision-language model and train it using navigation samples collected from continuous R2R datasets.
+However, most of them map navigation inputs to action decisions directly without the reasoning output.
 In contrast, we train the open-source LLM to generate self-improving embodied reasoning explicitly to improve action decision accuracy, which enhances the decision interpretability as well as mitigates overfitting to training reasoning labels.
-Embodied Chain-of-thoughts Training.
-Chain-of-thoughts (CoT) reasoning wei2022chain has been a widely utilized technique in Large Language Models (LLMs) and Vision-Language Models (VLMs).
-By generating the intermediate reasoning steps rather than directly predicting the answer, CoT can promote the answer accuracy for various tasks such as mathematical reasoning, commonsense reasoning, code generation, etc openai ; OpenAI_2023 ; liu2024deepseek .
-Inspired by this, some recent works have trained LLMs/VLMs to generate reasonable CoT for improving the action decision accuracy in embodied tasks zawalski2024robotic ; lin2024navcot ; mu2023embodiedgpt ; zhao2025cot ; liu2025spatialcot ; zhou2024navgpt .
-ECoT zawalski2024robotic trains vision-language-action models to generate embodied reasoning including object bounding boxes and end effector positions to encourage better adaptation for robotic manipulation tasks.
-NavGPT-2 zhou2024navgpt resorts to the GPT-4V model to collect free-form step-wise CoT reasoning data to improve the navigational reasoning ability of the VLN-specialized models Chen2022ThinkGA .
-CoT-VLA zhao2025cot incorporates explicit visual chain-of-thought reasoning into vision-language-action models by predicting future image frames before generating the action sequence.
-In our work, we propose a novel self-improving embodied CoT training paradigm for boosting the VLN, which effectively improves the reasoning ability and decision accuracy of the navigation model in various unseen scenarios.
-Moreover, we construct formalized CoT supervision, which can reduce redundant reasoning information and simultaneously promote the model’s inference speed.
-Figure 2: Overview of our EvolveNav.
-EvolveNav introduces a two-stage training framework for fulfilling self-improving embodied reasoning.
+2.3 Embodied Chain-of-thoughts Training
+Chain-of-thoughts (CoT) reasoning [56] has been a widely utilized technique in Large Language Models (LLMs) and Vision-Language Models (VLMs).
+By generating the intermediate reasoning steps rather than directly predicting the answer, CoT can promote the answer accuracy for various tasks such as mathematical reasoning, commonsense reasoning, code generation, etc [53, 28, 57].
+Inspired by this, some recent works have trained LLMs/VLMs to generate reasonable CoT for improving the action decision accuracy in embodied tasks [30, 32, 58, 31, 59, 43].
+ECoT [30] trains vision-language-action models to generate embodied reasoning including object bounding boxes and end effector positions to encourage better adaptation for robotic manipulation tasks.
+NavGPT-2 [43] resorts to the GPT-4V model to collect free-form step-wise CoT reasoning data to improve the navigational reasoning ability of the VLN-specialized models [15].
+CoT-VLA [31] incorporates explicit visual chain-of-thought reasoning into vision-language-action models by predicting future image frames before generating the action sequence.
+In this work, we propose a novel self-improving embodied CoT training paradigm for boosting VLN performance, which effectively improves the reasoning ability and decision accuracy of the navigation model in various unseen scenarios.
+Moreover, we construct CoT supervision in a formalized manner, which significantly reduces redundant reasoning information and simultaneously promotes the model’s inference speed.
+Figure 2: Overview of EvolveNav.
+EvolveNav involves a two-phase training framework for fulfilling self-improving embodied reasoning.
 In Stage 1 Formalized CoT Supervised Fine-Tuning, the navigation agent is trained using pre-constructed formalized CoT labels to generate navigational reasoning by predicting the landmark needed to locate with the corresponding direction.
@@
 3.2).
-Finally, we describe our proposed self-improving embodied reasoning framework in detail (Sec.
+Finally, we delve into the details of the proposed self-improving embodied reasoning framework (Sec.
 3.3).
 The overview of our EvolveNav is presented in Figure 2.
-Specifically, our EvolveNav consists of two stages: 1) Formalized CoT Supervised Fine-Tuning (Sec.
-3.3.1), where the LLM-based VLN model is initially trained using formalized CoT labels with supervised fine-tuning to activate the model’s navigational reasoning ability and simultaneously promote the reasoning speed; 2) Self-Reflective Post-Training (Sec.
+Specifically, EvolveNav consists of two training stages: 1) Formalized CoT Supervised Fine-Tuning (Sec.
+3.3.1), where we curate formalized CoT labels to initially train the LLM-based VLN model with supervised fine-tuning to activate the model’s navigational reasoning ability and simultaneously promote the reasoning speed; 2) Self-Reflective Post-Training (Sec.
 3.3.2), where the model is further trained with its own reasoning outputs as self-enriched CoT labels to increase supervision diversity, accompanied by a self-reflective auxiliary task to encourage better learning of accurate navigational reasoning patterns by discriminating from incorrect ones.
@@
 3.2 Model Architecture
-We build the LLM-based navigation agent that can simultaneously produce the navigational chain-of-thought reasoning and action prediction, modifying from a recent LLM-based VLN work, NaviLLM zheng2024towards .
-Concretely, the navigation agent consists of a scene encoder $F_{v}$, an LLM backbone $F_{\mathrm{LLM}}$, and an action prediction head $F_{\mathrm{action}}$.
+We build the LLM-based navigation agent that can simultaneously produce the navigational chain-of-thought reasoning and action prediction, modifying from a recent LLM-based VLN work, NaviLLM [45].
+The navigation agent consists of a scene encoder $F_{v}$, an LLM backbone $F_{\mathrm{LLM}}$, and an action prediction head $F_{\mathrm{action}}$.
 At each timestep $t$, the agent receives the navigation instruction $I$, panoramic observation $O_{t}$, and navigation history features $H_{t}=\{h_{0},...,h_{t-1}\}$.
@@
 The navigation prompt $P$ is then constructed by integrating the tokenized instruction, the visual representations $\{V_{t,n}\}_{n=1}^{N}$, and navigation history features $H_{t}$.
-As shown in Figure 2, special tokens <hist> and <cand> are introduced as placeholder tokens, where we insert the features $H_{t}$ and $\{V_{t,n}\}_{n=1}^{N}$, respectively.
-In contrast to NaviLLM zheng2024towards that directly maps the navigational inputs to action decision, in our EvolveNav, we construct the following output hint in the prompt $P$ to guide the navigation agent to generate both the action decision and explicit navigational reasoning: “-Action Decision: <cls>.
+As shown in Figure 2, special tokens <<hist>> and <<cand>> are introduced as placeholder tokens, where we insert the features $H_{t}$ and $\{V_{t,n}\}_{n=1}^{N}$, respectively.
+In contrast to NaviLLM [45] that directly maps the navigational inputs to action decision, in EvolveNav, we construct the following output hint in the prompt $P$ to guide the navigation agent to generate both the action decision and explicit navigational reasoning: “-Action Decision: <<cls>>.
 -Navigational Reasoning: ”.
-The <cls> token is also a special token for facilitating subsequent action predictions.
-The navigation prompt $P$ is fed into the LLM backbone $F_{\mathrm{LLM}}$ to obtain the feature ${f}^{cls}_{t}$ of the <cls> token and the chain-of-thought (CoT) reasoning $\mathrm{CoT}$:
+The <<cls>> token is also a special token for facilitating subsequent action predictions.
+The navigation prompt $P$ is fed into the LLM backbone $F_{\mathrm{LLM}}$ to obtain the feature ${f}^{cls}_{t}$ of the <<cls>> token and the chain-of-thought (CoT) reasoning $\mathrm{CoT}$:
 ${f}^{cls}_{t},\mathrm{CoT}$ $=F_{\mathrm{LLM}}(P).$ (2)
@@
 ${a}_{t}$ $=F_{\mathrm{action}}({f}^{cls}_{t}).$ (3)
-3.3 Self-Improving Embodied Reasoning
+3.3 Self-Improving Embodied Reasoning Framework
 3.3.1 Stage 1: Formalized CoT Supervised Fine-Tuning
@@
 When facing a given human instruction, the navigation agent usually needs to sequentially reason about the direction or the landmark it should move to in its current visual observation to reach the target position.
-Therefore, in our EvolveNav, we train the LLM-based VLN model to generate the chain-of-thought (CoT) reasoning about the landmark with the corresponding direction at different navigation timesteps, like the following format: “I should go to an observation with [landmark] to the [direction] of me”.
+Therefore, in EvolveNav, we train the LLM-based VLN model to generate the chain-of-thought (CoT) reasoning about the landmark with the corresponding direction at different navigation timesteps, like the following format: “I should go to an observation with [landmark] to the [direction] of me”.
 To encourage the navigation agent to choose the ground-truth action $a^{*}_{t}$ (paired observation is denoted as $O_{t}^{*}$) at different timesteps $t$ through the guidance of CoT reasoning during training, we obtain the corresponding landmarks $L$ and direction $D$ of $O_{t}^{*}$ to construct the formalized CoT labels, which is described as follows.
 Denote the observation $O_{t}^{*}$ as $O_{t}^{*}=\{B_{t},A_{t}=\{\psi_{t},\theta_{t}\}\}$, where $B_{t}$ is the RGB image of $O_{t}^{*}$, $A_{t}$ represent the direction information containing heading $\psi_{t}$ and elevation $\theta_{t}$.
-For the image $B_{t}$, we first employ a powerful image captioning model huang2023tag2text $F_{\mathrm{cap}}$ to obtain object and scene context $C_{t}$:
+For the image $B_{t}$, we first employ a powerful image captioning model [60] $F_{\mathrm{cap}}$ to obtain object and scene context $C_{t}$:
 $C_{t}=F_{\mathrm{cap}}(B_{t}).$ (4)
-Then, we leverage the NLP tool Spacy spacy to extract the landmarks list $L$ from $C_{t}$.
+Then, we leverage the NLP tool Spacy [61] to extract the landmarks list $L$ from $C_{t}$.
 In contrast to directly using object recognition models which may detect multiple redundant objects, extracting landmarks from the image captions can better retain salient landmarks.
 As a result, the generated CoT reasoning of the navigation agent can effectively help it locate important landmarks mentioned in the human instruction, since humans also tend to focus on salient landmarks when giving navigation instructions.
-We follow lin2024navcot to map the direction information $A_{t}$ of the observation $O_{t}^{*}$ to textual represented direction $D$.
+We follow [32] to map the direction information $A_{t}$ of the observation $O_{t}^{*}$ to textual represented direction $D$.
 With the landmarks list $L$ and direction $D$, we construct CoT labels $\mathrm{\mathrm{CoT}^{*}}$ by filling the following label template: I should go to an observation with [$L$] to the [$D$] of me.
-Through using our formalized CoT labels for training, the generation of redundant reasoning information can be effectively alleviated to promote the reasoning speed.
+Through extracting the landmark and direction information of the ground-truth observation (action) straightforwardly to construct the CoT labels, we do not explicitly correlate the CoT labels and the navigation instructions, leading to excellent generalization to navigational instructions of various types.
+Such CoT label construction strategy can effectively alleviate the problem that some action decisions are not explicitly corresponding to the navigation instruction, while instead enabling latent alignment learning of action decision, CoT reasoning, and navigational inputs.
 Supervised Fine-Tuning with Formalized CoT Labels.
@@
 $\mathcal{L}_{\mathrm{Stage1}}=\mathcal{L}_{\mathrm{action}}+\lambda\mathcal{L}_{\mathrm{SFT}},$ (6)
-where $\lambda$ represents the loss balance factor. We follow zheng2024towards to calculate the action prediction training objective $\mathcal{L}_{\mathrm{action}}$.
+where $\lambda$ represents the loss balance factor. We follow [45] to calculate the action prediction training objective $\mathcal{L}_{\mathrm{action}}$.
+Merits of formalized CoT labels.
+Our design of formalized CoT labels has the following merits compared with free-form CoT labels like those collected in [43] (also see Figure 2): Firstly, as the VLN task needs sequential decision making, generating formalized CoTs can significantly promote the reasoning speed compared to generating free-form ones.
+Secondly, free-form CoTs created by modern frontier models like GPT-4 [43] may produce irrelevant and redundant reasoning for decision, while formalized CoT can produce concise and task-related reasoning.
+Thirdly, using formalized CoT labels for training can effectively simplify the training process as well as mitigate the hallucination compared to using free-form ones.
+TABLE I: Performance comparison results on R2R under the task-specific training setting.
+* denote our reimplementation results.
+IL means the imitation learning setting.
+The best results for Cross-Modal Backbone and LLM-based Backbone are annotated in blue and bold fonts, respectively.
+Method Val Unseen Test Unseen
+[…]
+Cross-Modal Backbone:
+PREVALENT [26] 10.19 4.71 58 53 10.51 5.30 54 51
+HOP [62] 12.27 3.80 64 57 12.68 3.83 64 59
+HAMT [13] 11.46 2.29 66 61 12.27 3.93 65 60
+VLN-BERT [12] 12.01 3.93 63 57 12.35 4.09 63 57
+DUET [15] 13.94 3.31 72 60 14.73 3.65 69 59
+Meta-Explore [63] 13.09 3.22 72 62 14.25 3.57 71 61
+VLN-SIG [64] - - 72 62 - - 72 60
+VLN-PETL [65] 11.52 3.53 65 60 12.30 4.10 63 58
+NavGPT2 [43] 13.25 3.18 71 60 - - - -
+LLM-based Backbone:
+NavGPT [24] 11.45 6.46 34 29 - - - -
+DiscussNav [25] 9.69 5.32 43 40 - - - -
+MapGPT [23] - 5.63 34 29 - - - -
+NavCoT [32] 9.95 6.26 40 37 - - - -
+NaviLLM* [45] (IL) 9.99 6.04 46.90 43.78 10.03 6.12 46 43
+EvolveNav (IL, ours) 9.79 5.52 51.15 48.27 9.94 5.92 47 45
+NaviLLM* [45] 13.43 3.27 70.11 60.25 13.68 3.37 70 61
+EvolveNav (ours) 12.07 3.15 71.17 63.48 12.06 3.22 71 63
 3.3.2 Stage 2: Self-Reflective Post-Training
 Although CoT supervised fine-tuning can explicitly guide the navigation agent to produce navigational reasoning for assisting the action decision, due to the uncertainty and complexity of the navigation task, using fixed labels may lead to overfitting to training CoT label distributions and therefore harm the generalization to unseen scenarios.
-Moreover, the inherent noise in the image captioning model huang2023tag2text for landmark detection may also limit the accuracy of formalized CoT labels collected in Stage 1.
-Therefore, after the Stage 1 training of Formalized CoT Supervised Fine-Tuning, we introduce Self-Reflective Post-Training to further encourage the navigation agent to learn correct reasoning patterns for improving generalization.
+Moreover, the inherent noise in the image captioning model [60] for landmark detection may also limit the accuracy of formalized CoT labels collected in Stage 1.
+Therefore, after the Stage 1 training of Formalized CoT Supervised Fine-Tuning, we introduce Self-Reflective Post-Training to further encourage the navigation agent to learn correct reasoning patterns in a self-improving manner for improving generalization.
 Self-Enriched CoT Label Enhancement.
@@
 We utilize the above mentioned CoT label $\tilde{\mathrm{CoT}}_{t}^{*}$ as the positive reasoning sample $R^{+}$.
-To obtain the negative reasoning sample $R^{-}$, we randomly select the candidate observation (action) $O_{t,j}$ ($1<j<N$, $N$ is the number of navigable views) different from the ground-truth one.
+To obtain the negative reasoning sample $R^{-}$, we randomly select the candidate observation (action) $O_{t,j}$ ($1<j<N$, NN is the number of navigable views) different from the ground-truth one.
 Then we extract the landmark and direction for $O_{t,j}$ to fill in the CoT label template (see Sec.
@@
 As shown in Figure 2, we construct the self-reflective task prompt $P_{\mathrm{sr}}$ as “Choose the correct one from the given two navigational reasoning outputs.
-Output 1: [$R^{1}$].
-Output 2: [$R^{2}$].
+Output 1: $R^{1}$].
+Output 2: $R^{2}$].
 Selection: ”, where we randomly insert the positive reasoning sample $R^{+}$ and negative reasoning sample $R^{-}$ to the positions of $R^{1}$ and $R^{2}$.
@@
 Through the self-enriched CoT label enhancement and self-reflective auxiliary task, the navigation agent learns to generate correct embodied reasoning in a self-refining manner to enable adaptable reasoning in different scenarios.
+3.3.3 Training and Inference
+Two-stage training.
+When conducting training in the proposed self-improving embodied reasoning framework, we train the model in Stage 1 to converge, and use it for Stage 2 training.
+This two-stage training design is to mitigate the negative impact of noisy CoT outputs during Stage 1 under a non-converged situation and ensure the training stability during Stage 2.
+Specifically, in the early training iterations during Stage 1 of Formalized CoT Supervised Fine-Tuning, the agent is prone to generate CoT reasoning with noisy formats and information (e.g., its output may contain repeatedly notions of “<<s>>” or the output may be very long) while accompanied action decisions may be correct occasionally.
+Based on the rule of our self-enriched label enhancement strategy during Stage 2, such noisy outputs will be introduced as CoT labels to guide the CoT training and may cause training instability.
+By firstly training the model in Stage 1 to converge, such issues can be effectively alleviated.
+This two-stage training manner also simplifies the method design and implementation to promote its practicality during real deployment.
+Inference.
+During inference, the model generates CoT based on the given prompt with the form of “-Action Decision: <<cls>>.
+-Navigational Reasoning:”.
+The encoding of the <<cls>> token is extracted to generate the action prediction probability $a_{t}$ through the action prediction head $F_{\mathrm{action}}$ (see Sec.
+3.2).
+TABLE II: Performance comparison results on CVDN under the task-specific training setting.
+We utilize the Goal Progress (GP) (m) as the evaluation metric.
+* denote our reimplementation results.
+The best results for Cross-Modal Backbone and LLM-based Backbone are annotated in blue and bold fonts, respectively.
+Method Val-Unseen Test
+Cross-Modal Backbone:
+Seq2Seq [34] 2.10 2.35
+PREVALENT [26] 3.15 2.44
+HOP [62] 4.41 3.31
+MT-RCM [66] 4.36 -
+MT-RCM+Env [66] 4.65 3.91
+HAMT [13] 5.13 5.58
+VLN-SIG [64] 5.52 5.83
+VLN-PETL [65] 5.69 6.13
+LLM-based Backbone:
+NaviLLM* [45] 5.53 6.80
+EvolveNav (ours) 6.21 7.07
 4 Experiment
 4.1 Experimental Setup
-Datasets.
-We test our EvolveNav on two popular VLN benchmarks, i.e., R2R anderson2018vision and CVDN thomason2019vision .
+4.1.1 Datasets
+We test EvolveNav on four popular VLN benchmarks, i.e., R2R [33], CVDN [34], REVERIE [35], and SOON [36].
+Each benchmark handles distinct challenges posed by VLN.
 R2R is built on 90 real-world indoor simulation environments containing 7,189 trajectories, each corresponding to three fine-grained instructions.
 CVDN contains 2,050 human-human navigation dialogs and over 7k trajectories in 83 MatterPort houses.
-Evaluation Metrics.
-We utilize the following standard metrics for evaluation: 1) Trajectory Length (TL): the average length of the agent’s navigated path, 2) Navigation Error (NE): the average distance between the agent’s destination and the goal viewpoint, 3) Success Rate (SR): the ratio of success, where the agent stops within three meters of the target point, 4) Success rate weighted by Path Length (SPLanderson2018vision : success rate normalized by the ratio between the length of the shortest path and the predicted path, 5) Oracle Success Rate (OSR): the ratio of containing a viewpoint along the path where the target position is visible, 6) Goal Progress (GP), the progress in meters towards the goal.
-Table 1: Comparison with state-of-the-art methods on the R2R dataset. * denotes training with R2R in-domain data. IL means the imitation learning setting.
-Method Val Unseen Test Unseen
+REVERIE replaces the fine-grained instructions in R2R with high-level instructions.
+SOON constructs thoroughly described instructions to further highlight visual-semantic alignment.
+To verify the effectiveness of EvolveNav, we adopt two representative training setting, task-specific training and cross-task training.
+Task-specific training trains the model on single benchmark like most previous works [13, 15, 32], while cross-task training realizes a generalist navigation model like NaviLLM [45] by training the model using multiple datasets.
+4.1.2 Evaluation Metrics
+We utilize the following standard metrics for evaluation: 1) Trajectory Length (TL): the average length of the agent’s navigated path, 2) Navigation Error (NE): the average distance between the agent’s destination and the goal viewpoint, 3) Success Rate (SR): the ratio of success, where the agent stops within three meters of the target point, 4) Success rate weighted by Path Length (SPL[33]: success rate normalized by the ratio between the length of the shortest path and the predicted path, 5) Oracle Success Rate (OSR): the ratio of containing a viewpoint along the path where the target position is visible, 6) Goal Progress (GP), the progress in meters towards the goal.
+4.1.3 Implementation Details
+We utilize NaviLLM [45] as our baseline model.
+To simplify the implementation, we do not introduce the pretraining phase in [45] for both our EvolveNav and the baseline (denoted as NaviLLM* in Table IV).
+In our EvolveNav, We fine-tune the LLM with full-parameter and LoRA settings for Stage 1 and 2 training, respectively.
+The training for Stage 1 is conducted on 8 Nvidia A100 GPUs and the training for Stage 2 is performed on 4 Nvidia A100 GPUs.
+Empirically, we set the loss coefficients $\lambda$, $\lambda_{1}$, and $\lambda_{2}$ as 1, 1, and 0.2, respectively.
+During Stage 1 training, we introduce the CoT supervised finetuning loss $\mathcal{L}_{\mathrm{SFT}}$ under a probability of 0.5 to mitigate the overfitting to the pre-constructed CoT labels.
+The maximum numbers of training steps for Stage 1 and 2 are set as 60000 and 9000 steps, respectively.
+Training for Stage 1 with full parameter lasts for $\sim$1.5 days with $\sim$73G GPU memory, and training for Stage 2 with LoRA lasts for $\sim$1 day with $\sim$30G GPU memory.
+The hyperparameters such as the learning rate, optimizer, and the batch size are kept the same as [45].
+TABLE III: Performance comparison results on SOON under the task-specific training setting.
+* denote our reimplementation results.
+The best results for Cross-Modal Backbone and LLM-based Backbone are annotated in blue and bold fonts, respectively.
+Method Val-Unseen
 […]
 Cross-Modal Backbone:
-PREVALENT hao2020towards 10.19 4.71 58 53 10.51 5.30 54 51
-HOP qiao2022hop 12.27 3.80 64 57 12.68 3.83 64 59
-HAMT Chen2021HistoryAM 11.46 2.29 66 61 12.27 3.93 65 60
-VLN-BERT hong2021vln 12.01 3.93 63 57 12.35 4.09 63 57
-DUET Chen2022ThinkGA 13.94 3.31 72 60 14.73 3.65 69 59
-Meta-Explore Hwang2023MetaExploreEH 13.09 3.22 72 62 14.25 3.57 71 61
-VLN-SIG li2023improving - - 72 62 - - 72 60
-VLN-PETL qiao2023vln 11.52 3.53 65 60 12.30 4.10 63 58
-NavGPT2 zhou2024navgpt 13.25 3.18 71 60 - - - -
+GBE [36] 28.54 19.52 13.34
+DUET [15] 50.91 36.28 22.58
+AZHP [67] 56.19 40.71 26.58
 LLM-based Backbone:
-NavGPT zhou2023navgpt 11.45 6.46 34 29 - - - -
-DiscussNav long2023discuss 9.69 5.32 43 40 - - - -
-MapGPT chen2024mapgpt - 5.63 34 29 - - - -
-NavCoT lin2024navcot 9.95 6.26 40 37 - - - -
-NaviLLM zheng2024towards 12.81 3.51 67 59 13.21 3.71 68 60
-NaviLLM* (IL) 9.99 6.04 46.90 43.78 10.03 6.12 46 43
-EvolveNav (IL) 9.79 5.52 51.15 48.27 9.94 5.92 47 45
-NaviLLM* 13.43 3.27 70.11 60.25 13.68 3.37 70 61
-EvolveNav 12.07 3.15 71.17 63.48 12.06 3.22 71 63
-Implementation Details.
-We fine-tune the LLM with full-parameter and LoRA settings for Stage 1 and 2, respectively.
-To simplify the implementation, we do not introduce the pretraining phase like zheng2024towards for both our EvolveNav and the compared baseline (denoted as NaviLLM* in Table 1 &  2).
-The training for Stage 1 is conducted on 8 Nvidia A100 GPUs and the training for Stage 2 is performed on 4 Nvidia A100 GPUs.
-Empirically, we set the loss coefficients $\lambda$, $\lambda_{1}$, and $\lambda_{2}$ as 1, 1, and 0.2, respectively.
-The hyperparameters such as the learning rate, optimizer, and the batch size are kept the same as zheng2024towards .More implementation details are presented in the supplementary material.
+NaviLLM* [45] 43.60 30.34 23.70
+EvolveNav (Ours) 49.56 33.40 24.92
+TABLE IV: Performance comparison results on REVERIE under the task-specific training setting.
+* denote our reimplementation results.
+The best results for Cross-Modal Backbone and LLM-based Backbone are annotated in blue and bold fonts, respectively.
+Method Val Unseen
+[…]
+Cross-Modal Backbone:
+Seq2Seq [33] 8.07 4.20 2.84
+HOP [62] 36.24 31.78 26.11
+HAMT [13] 36.84 32.95 30.20
+VLN-BERT [12] 35.02 30.67 24.90
+DUET [15] 51.07 46.98 33.73
+AZHP [67] 53.65 48.31 36.63
+VLN-PETL [65] 37.03 31.81 27.67
+LLM-based Backbone:
+NaviLLM* [45] 42.68 32.55 25.82
+EvolveNav (Ours) 42.40 33.60 28.16
 4.2 Comparison with Existing Methods
-Table 1 shows the comparison among different approaches on R2R.
-We compare EvolveNav with the baseline method NaviLLM zheng2024towards using the R2R in-domain training data (denoted as NaviLLM*) under two settings, i.e., imitation learning (IL) and Dagger training zheng2024towards ; Chen2022ThinkGA .
-The comparison shows the consistent improvement of our approach over NaviLLM* on both Val Unseen and Test Unseen under different training settings.
-For example, the performance gain in SPL of EvolveNav on Val Unseen is $\sim$4.5and $\sim$3.2under IL and dagger training settings compared to the baseline, respectively.
-Note that our EvolveNav achieves the state-of-the-art performance on R2R among the approaches with LLM-based backbone.
-These results demonstrate the effectiveness of our self-improving embodied reasoning framework in fulfilling generalizable navigational reasoning.
-We do not consider the comparison with models augmented by new environments (e.g., ScaleVLN wang2023scaling ) for fairness.
-Table 2: Comparison with SoTA methods on CVDN. We utilize the Goal Progress (GP) (m) as the evaluation metric. * denotes training with CVDN in-domain data.
-Method Val-Unseen Test
-Cross-Modal Backbone:
-Seq2Seq thomason2019vision 2.10 2.35
-PREVALENT hao2020towards 3.15 2.44
-HOP qiao2022hop 4.41 3.31
-MT-RCM wang2020environment 4.36 -
-MT-RCM+Env wang2020environment 4.65 3.91
-HAMT Chen2021HistoryAM 5.13 5.58
-VLN-SIG li2023improving 5.52 5.83
-VLN-PETL qiao2023vln 5.69 6.13
-LLM-based Backbone:
-NaviLLM zheng2024towards 6.16 7.90
-NaviLLM* 5.53 6.80
-EvolveNav (ours) 6.21 7.07
-In Table 2, we compare our EvolveNav with different VLN approaches on CVDN.
-As shown in Table 2, EvolveNav outperforms the baseline NaviLLM* zheng2024towards on both Val Unseen and Test set on CVDN, demonstrating good generalization ability of our proposed self-improving embodied reasoning paradigm to different kinds of VLN tasks.
-For example, EvolveNav brings $\sim$0.7 performance gain over NaviLLM* in GP on Val Unseen.
-Moreover, Table 2 shows the superiority of EvolveNav over NaviLLM zheng2024towards , which conducts multi-task training on multiple VLN benchmarks to learn cross-task navigation knowledge, demonstrating the importance of explicit navigational reasoning learning compared to direct input-output mapping paradigms.
-Such results also reveal the effectiveness of our constructed CoT labels, which provide landmark and direction information for enabling the agent to learn useful navigational reasoning for decisions.
-The results on both R2R and CVDN benchmarks demonstrate the effectiveness of our EvolveNav in fulfilling generalizable navigational reasoning to improve navigation decision accuracy.
+Task-specific training.
+Table IIIIII, and IV present the task-specific training results on R2R [33], CVDN [34], SOON [36], and REVERIE [35], respectively.
+The results exhibit consistent superiority of EvolveNav over the compared approaches on various VLN benchmarks, demonstrating the effectiveness and excellent generalization ability of the proposed self-improving embodied reasoning paradigm.
+For example, for the results on R2R in Table I, the performance gain in SPL of EvolveNav on Val Unseen is $\sim$4.5% and $\sim$3.2% under the imitation learning (IL) and dagger [15] training settings compared to the baseline model NaviLLM [45], respectively.
+For the results on SOON in Table [36], the performance improvements of EvolveNav over the baseline model NaviLLM on OSR, SR, and SPL are $\sim$5.9%, $\sim$3.1%, and $\sim$1.2%, respectively.
+Note that we do not consider the comparison with models augmented by new environments (e.g., ScaleVLN [41]) for fairness.
+TABLE V: Performance comparison results under the cross-task training setting.
+* denote our reimplementation results.
+The best results for task-specific training (“Separate Model for Each Task”) and cross-task training (“Unified Model for All Tasks”) are annotated in blue and bold fonts, respectively.
+Method REVERIE SOON R2R CVDN
+[…]
+Separate Model for Each Task:
+DUET [15] 46.98 51.07 33.73 36.28 50.91 22.58 72 - 60 -
+AZHP [67] 48.31 53.65 36.63 40.71 56.19 26.58 - - - -
+VLN-PETL [65] 31.81 37.03 27.67 - - - 65 - 60 5.69
+NaviLLM* [45] 32.55 42.68 25.82 30.34 43.60 23.70 70.11 79.00 60.25 5.53
+EvolveNav (Ours) 33.60 42.40 28.16 33.40 49.56 24.92 71.17 78.95 63.48 6.21
+Unified Model for All Tasks:
+NaviLLM [45] 44.56 53.74 36.63 35.44 - 28.09 67 - 58 5.91
+NaviLLM* [45] 43.81 54.26 35.61 34.82 55.01 26.19 68.37 76.96 59.09 5.43
+EvolveNav (Ours) 45.97 57.95 38.58 37.00 58.20 28.18 68.07 81.68 58.30 6.35
+Cross-task training.
+Table V shows the cross-task training results on R2R [33], CVDN [34], SOON [36], and REVERIE [35].
+From Table V, we can observe that EvolveNav surpasses the baseline approach NaviLLM [45] in most metrics on different benchmarks.
+These results reveal that our self-improving embodied reasoning framework is also effective for training the generalist navigation model, which is more practical and flexible in real-world navigation scenarios.
+Both the task-specific and cross-task training results on various VLN benchmarks sufficiently demonstrate that the proposed self-improving embodied reasoning framework fulfills adaptable and generalizable navigational reasoning under different tasks and scenarios.
 4.3 Ablation Study
 Effect of different method components.
-Table 3 presents the ablation study on Val Unseen set on R2R, where we can find the effectiveness and reasonability of different method components in our EvolveNav.
-From Table 3, we can observe that through the Stage 1 training of formalized CoT supervised fine-tuning (SFT), the model’s navigational reasoning ability can be significantly enhanced to improve the navigation performance.
-The introduction of self-enriched CoT labels and the self-reflective auxiliary task can further bring performance gain in both SR and SPL metrics compared to pure CoT SFT.
-Our full model achieves the best results in SR and SPL among all variants, demonstrating the effectiveness of our two-phase self-improving embodied training framework.
-Table 3: Ablation study of method components on Val Unseen set on R2R. We adopt the imitation learning (IL) setting for evaluation.
+Table VI presents the ablation study results on Val Unseen set on R2R, where we can find the effectiveness and reasonability of different method components in our EvolveNav.
+From Table VI, we can observe that through the Stage 1 training of formalized CoT supervised fine-tuning (SFT) (“1”), the model’s navigational reasoning ability can be significantly enhanced to improve the navigation performance.
+In Stage 2 training, the introduction of self-enriched CoT labels (“2”) and the self-reflective auxiliary task (“3”) can further bring performance gain respectively in both SR and SPL metrics compared to pure CoT SFT in Stage 1.
+Our full model (“Full Model”) achieves the best results in SR and SPL compared with “2” and “3”, demonstrating that the combination of our self-enriched CoT label enhancement strategy and self-reflective auxiliary task is non-trivial.
+Especially, rather than pure self-reflective auxiliary task (“3”) that asks the agent to learn to distinguish fixed correct and wrong reasoning patterns like conventional auxiliary task design, the combination of our self-enriched CoT label enhancement strategy and self-reflective auxiliary task (“Full Model”) can encourage the agent to learn diverse correct reasoning and therefore increase its generalization ability to unseen scenarios.
+TABLE VI: Ablation study of method components on Val Unseen set on R2R.
+We adopt the imitation learning (IL) setting for evaluation.
+“CoT SFT” represents the Stage 1 training of Formalized CoT Supervised-Finetuning.
+“Self-Enriched CoT SFT” denote the Self-Enriched CoT Label Enhancement strategy in Stage 2 training of Self-Reflective Post-Training.
+Method Stage 1 Stage 2 Val Unseen
+CoT
+SFT
+Self-Enriched
+CoT SFT
+Self-Reflective
+Auxiliary Task
 […]
-Baseline 46.90 54.63 43.78
-   + Formalized CoT SFT 49.62 59.44 46.26
-   + Self-Enriched CoT SFT 50.47 57.48 47.98
-   + Self-reflective task 50.51 57.74 47.86
-   + Full Model (ours) 51.15 59.18 48.27
+Baseline - - - 46.90 54.63 43.78
+1 49.62 59.44 46.26
+2 50.47 57.48 47.98
+3 50.51 57.74 47.86
+Full Model 51.15 59.18 48.27
+Figure 3: Action decision visualization of NaviLLM [45] and our EvolveNav.
+We only extract two steps and display local candidate space for simplicity.
+Observations selected by EvolveNav (also are the ground-truth actions) and NaviLLM are annotated by green boxes and red boxes, respectively.
 Effect of constructed CoT labels.
-Table 4 compares the navigation performance under different kinds of CoT labels, where we can find the effectiveness of our introduced CoT labels by predicting landmark and direction information in a formalized way.
-To realize “Free-form CoT”, we introduce the free-form CoT labels collected in NavGPT-2 zhou2024navgpt to train the navigation agent.
-To realize “Direction & Landmark”, we obtain the best matched landmark in the instruction to each ground-truth observation through the CLIP model radford2021learning .
-The comparison between “Free-form CoT” and “Direction & Landmark (ours)” shows that our formalized CoT labels can effectively reduce redundant reasoning information to improve the navigational reasoning and decision accuracy.
-Moreover, the inference time to generate CoT at one timestep of “Free-form CoT” is $\sim$7.8s compared to $\sim$2.5s of “Direction & Landmark (ours)”, demonstrating that our method can significantly promote the reasoning speed ($\sim$×$\times$3 improvement), which is crucial for sequential decision making task like navigation.
-By comparing “Only Direction” and “Only Landmark” with “Direction & Landmark (ours)”, we can find that both landmark and direction information are important to guide navigation decisions, revealing the reasonability of our constructed CoT labels.
-The superiority of “Direction & Landmark (ours)” over “Direction & Landmark” demonstrates that our introduced CoT labels, which contain diverse landmarks, can potentially encourage the navigation agent to learn cross-modal alignment knowledge to accurately align the visual observation to the navigation instruction.
-Table 4: Ablation study of CoT on Val Unseen set on R2R. denotes using the landmark mentioned in the instruction.
+Table VII compares the navigation performance under different kinds of CoT labels, where we can find the effectiveness of our introduced CoT labels by predicting landmark and direction information in a formalized way.
+To realize “Free-form CoT”, we introduce the free-form CoT labels collected in NavGPT-2 [43] to train the navigation agent.
+To realize “Direction & Landmark”, we obtain the best matched landmark in the instruction to each ground-truth observation through the CLIP model [68].
+From Table VII, we can draw into the following conclusions: 1) The comparison between “Free-form CoT” and “Direction & Landmark (ours)” shows that our formalized CoT labels can effectively reduce redundant reasoning information to improve the navigational reasoning and decision accuracy.
+Moreover, the inference time to generate CoT at one timestep of “Free-form CoT” is $\sim$7.8s compared to $\sim$2.5s of “Direction & Landmark (ours)”, demonstrating that our method can significantly promote the reasoning speed ($\sim$$\times$3 improvement), which is crucial for sequential decision making task like navigation.
+2) The comparison among “Only Direction”, “Only Landmark”, and “Direction & Landmark (ours)” reveal that both landmark and direction information are important to guide navigation decisions, demonstrating the reasonability of our constructed CoT labels.
+3) The superiority of “Direction & Landmark (ours)” over “Direction & Landmark” demonstrates that our introduced CoT labels, which contain diverse landmarks, can potentially encourage the navigation agent to learn cross-modal alignment knowledge to accurately align the visual observation to the navigation instruction.
+Figure 4: Visualization comparison between self-enriched chain-of-thought (CoT) labels and originally built CoT labels.
+Newly introduced landmarks in the self-enriched CoT label are highlighted in red fonts.
+GT action denotes the ground-truth action (observation).
+We omit the direction information in the CoT labels.
+TABLE VII: Ablation study of CoT labels on Val Unseen set on R2R. denotes using the landmark mentioned in the instruction.
 […]
@@
 4.4 Visualization
+In this subsection, we present various kinds of visualization results, including visualization of action decision, self-enriched CoT label, loss & performance variation, and landmark extraction, to comprehensively and deeply analyze the advantage of the proposed self-improving embodied reasoning framework.
+Action Decision Visualization.
 Fig.
-3 gives the action decision visualization comparison between NaviLLM zheng2024towards and our EvolveNav, where we can find that EvolveNav generates reasonable navigational reasoning about landmarks and directions to guide correct action decision making.
+3 gives the action decision visualization comparison between NaviLLM [45] and our EvolveNav, where we can find that EvolveNav generates reasonable navigational reasoning about landmarks and directions to guide correct action decision making.
 For example, when $t=0$, from the observations, EvolveNav infers that an observation with door and hallway represents the exit from the bedroom while NaviLLM mistakenly selects an action that remains in the bedroom.
@@
 These results highlight the effectiveness of our approach in improving navigational reasoning for accurate instruction understanding and action prediction.
-Figure 3: Action decision visualization of NaviLLM zheng2024towards and our EvolveNav.
-We only extract two steps and display local candidate space for simplicity.
-Observations selected by EvolveNav (also are the ground-truth actions) and NaviLLM are annotated by green boxes and red boxes, respectively.
-5 Conclusion
-In this work, we propose EvolveNav, a novel self-improving embodied reasoning framework to fulfill generalizable and adaptable reasoning for enhancing LLM-based vision-and-language navigation.
-Through introducing the formalized CoT supervised fine-tuning and self-reflective post-training in the proposed framework, the agent’s navigational reasoning ability can be effectively enhanced while mitigating the overfitting to the training reasoning label distributions simultaneously to improve generalization.
-Experimental results on two popular benchmarks R2R and CVDN reveal the promising capability of our method in boosting the reasoning ability and decision accuracy for LLM-based navigation agents.
-We believe that our EvolveNav can provide meaningful references for designing self-improving embodied reasoning paradigms to benefit future LLM-assisted Embodied AI research.
-Supplementary Material
-Appendix A Overview
-In these supplementary materials, we first analyze the broader impacts and limitations of our EvolveNav in Sec.
-B and Sec.
-C, respectively.
-Then, we describe more implementation details in Sec.
-D.
-Finally, we present more experimental results, including the landmark extraction visualization, self-enriched CoT label visualization, and loss & performance curves during Self-Reflective Post-Training, in Sec.
-E.
-Appendix B Broader Impact
-The scalable deployment of LLM-based navigation robots holds significant potential to enhance operational efficiency in public spaces, reduce human labor costs, and benefit diverse populations including the elderly and individuals with disabilities.
-Our proposed self-improving embodied reasoning framework can help improve both navigation success rate and decision interpretability of LLM-based navigation agents, therefore contributing to scalable application of LLM-based navigation robots which have high robustness and reliability.
-However, the inherent hallucination of LLMs may cause incorrect decisions in complex environments, leading to potential risks such as collisions.
-More efforts devoted to hallucination suppression of LLMs and safety promotion of navigation robots are encouraged for future research in the community.
-Appendix C Limitations
-(1) In the Stage 1 training of Formalized CoT Supervised Fine-tuning in our EvolveNav, we conduct full-parameter training for the open-source LLM, vicuna-7B [1], like the baseline method [66].
-This process costs $\sim$1.5 day training on 8 NVIDIA A100 GPUs with 80G memory, which is relative resource-intensive compared to conventional VLN approaches [10, 9, 17].
-In future work, we plan to introduce parameter-efficient training strategies or resort to more lightweight LLMs to reduce the computational resource consumption.
-(2) During the formalized CoT label collection for supervised fine-tuning, we employ the open-vocabulary image captioning model [24] to obtain the caption for visual observations, followed by the NLP tool Spacy [20] to extract the landmark from the caption.
-This process is performed in an offline manner, which might introduce noise for the CoT label construction due to the possible inaccuracy of the image captioning model and the NLP tool.
-Future work includes utilizing more powerful open-vocabulary models or conducting human checking to promote the accuracy of the CoT labels.
-Appendix D Implementation Details
-We adopt Vicuna 7B [1] as our large language model (LLM) backbone, following [66].
-For R2R [5], we also introduce the auxiliary tasks of trajectory summarization and embodied question answering in [66] during training.
-The maximum number of landmark extracted for constructing formalized CoT labels is set as 5 to retain useful information and simultaneously promote the reasoning speed.
-To mitigate the possible landmark missing during landmark detection due to the view deviation for a single candidate observation, we fuse the landmark information of candidate observations with the same direction to construct the CoT labels.
-We remove some uninformative landmarks such as “floor” and “ceiling” empirically during the landmark extraction.
-During Stage 1 training, we introduce the CoT supervised finetuning loss $\mathcal{L}_{\mathrm{SFT}}$ under a probability of 0.5 to mitigate the overfitting to the pre-constructed CoT labels.
-The maximum numbers of training steps for Stage 1 and 2 are set as 60000 and 9000 steps, respectively.
-Training for Stage 1 with full parameter lasts for $\sim$1.5 days with $\sim$73G GPU memory, and training for Stage 2 with LoRA lasts for $\sim$1 day with $\sim$30G GPU memory.
-The inference is conducted on 4 V100 GPUs with the evaluation batch size of 2.
-During inference, we set the LLM temperature to 0 and opt for a greedy strategy to generate action decisions.
-Appendix E More Experimental Results
+(a) Loss variation
+(b) Navigation Error (NE) variation
+Figure 5: Loss and performance variation during Stage 2: Self-Reflective Post-Training. Low navigation error (NE) value indicates better results.
+Figure 6: Landmark extraction visualization of different methods. We use green, red, and blue colors to distinguish informative, false, and uninformative landmarks, respectively.
+Self-Enriched CoT label visualization.
+In Fig.
+4, we present some visualization comparison examples of original CoT labels and self-enriched CoT labels.
+From Fig.
+4, we can observe that our self-enriched CoT label enhancement strategy effectively increases the supervision diversity.
+For example, in Fig.
+4(a) and (b), the self-enriched CoT labels from the model’s own reasoning outputs capture the landmark painting and chair which are not contained in the original CoT label, respectively.
+From Fig.
+4(c) and (d), we can find that the LLM-based navigation agent can also recognize the attribute of the landmark, e.g., it recognizes wood paneling and arched window, which indicate the material and shape of the landmark, respectively.
+Such CoT labels can help navigation agent learn to follow more fine-grained instructions.
+Benefiting from our self-enriched CoT label, the LLM-based navigation agent can reduce the overfitting to the original CoT label distributions and learn more diverse cross-modal alignment knowledge, and therefore promote the generalization to unseen scenarios.
+Loss & performance variation during Stage 2 training.
+Fig.
+5 shows the loss and performance curves during Self-Reflective Post-Training (Stage 2).
+In Fig.
+5, we can find that our self-reflective loss $\mathcal{L}_{\mathrm{sr}}$, the total training loss $\mathcal{L}_{\mathrm{Stage2}}$, and the navigation error (NE) have similar variation trends.
+Especially, both two loss curves and the NE curves achieve the lowest value around epoch 9.
+These results show the effectiveness of our constructed self-reflective auxiliary task during Self-Reflective Post-Training in improving the navigational reasoning and decision accuracy of the LLM-based navigation agent.
 Landmark extraction visualization.
@@
 From Fig.
-4, we can observe that Tag2Text can capture more informative landmarks while having less redundancy and illusion.
+6, we can observe that Tag2Text can capture more informative landmarks while having less redundancy and illusion.
 For example, for “Cand 2”, Tag2Text correctly detects a kitchen and living room, a dining table, and chair, while BLIP-v2 only detects a large dining table.
@@
 These results show that our combination of Tag2text model [60] and NLP tool for landmark extraction can effectively retain informative landmarks while reducing redundancy and illusion for constructing CoT labels, which can enable the agent to better learn cross-modal alignment between observations and instructions.
-Figure 4: Landmark extraction visualization of different methods. We use green, red, and blue colors to distinguish informative, false, and uninformative landmarks, respectively.
-Self-Enriched CoT label visualization.
-In Fig.
-5, we present some visualization comparison examples of original CoT labels and self-enriched CoT labels.
-From Fig.
-5, we can observe that our self-enriched CoT label enhancement strategy effectively increases the supervision diversity.
-For example, in Fig.
-5(a) and (b), the self-enriched CoT labels from the model’s own reasoning outputs capture the landmark painting and chair which are not contained in the original CoT label, respectively.
-From Fig.
-5(c) and (d), we can find that the LLM-based navigation agent can also recognize the attribute of the landmark, e.g., it recognizes wood paneling and arched window, which indicate the material and shape of the landmark, respectively.
-Such CoT labels can help navigation agent learn to follow more fine-grained instructions.
-Benefiting from our self-enriched CoT label, the LLM-based navigation agent can reduce the overfitting to the original CoT label distributions and learn more diverse cross-modal alignment knowledge, and therefore promote the generalization to unseen scenarios.
-Figure 5: Self-enriched chain-of-thought (CoT) label visualization.
-Newly introduced landmarks in the self-enriched CoT label are highlighted in red fonts.
-GT action denotes the ground-truth action (observation).
-We omit the direction information in the CoT label.
-(a) Loss variation
-(b) Navigation Error (NE) variation
-Figure 6: Loss and performance variation during Stage 2: Self-Reflective Post-Training. Low navigation error (NE) value indicates better results.
-Loss & performance variation during Stage 2 training.
-Fig.
-6 shows the loss and performance curves during Self-Reflective Post-Training (Stage 2).
-In Fig.
-6, we can find that our self-reflective loss $\mathcal{L}_{\mathrm{sr}}$, the total training loss $\mathcal{L}_{\mathrm{Stage2}}$, and the navigation error (NE) have similar variation trends.
-Especially, both two loss curves and the NE curves achieve the lowest value around epoch 9.
-These results show the effect of our self-reflective auxiliary task during Self-Reflective Post-Training in improving the navigational reasoning and decision accuracy of the LLM-based navigation agent.
+5 Conclusion
+In this work, we propose EvolveNav, a novel self-improving embodied reasoning framework to fulfill generalizable and adaptable reasoning for enhancing LLM-based vision-and-language navigation.
+Through introducing the formalized CoT supervised fine-tuning and self-reflective post-training in the proposed framework, the agent’s navigational reasoning ability can be effectively enhanced while mitigating the overfitting to the training reasoning label distributions simultaneously to improve generalization.
+Experimental results on multiple VLN benchmarks under diverse training settings reveal the promising capability of our method in boosting the reasoning ability and decision accuracy for LLM-based navigation agents.
+We believe that our EvolveNav can provide meaningful references for designing self-improving embodied reasoning paradigms to benefit future LLM-assisted Embodied AI research.